### v1

# Jamming is a first-order transition with quenched disorder in amorphous materials sheared by cyclic quasistatic deformations

Abstract
Jamming is an athermal transition between flowing and rigid states in amorphous systems such as granular matter, colloidal suspensions, complex fluids and cells. The jamming transition seems to display mixed aspects of a first-order transition, evidenced by a discontinuity in the coordination number, and a second-order transition, indicated by power-law scalings and diverging lengths. Here we demonstrate that jamming is a first-order transition with quenched disorder in cyclically sheared systems with quasistatic deformations. Particle models are simulated in two and three dimensions, which undergo reversible-irreversible transitions under cyclic shear. The ensemble of configurations, with partially crystallized and fragile states excluded, is generated in the irreversible phase, where the system is stationary and particles are diffusive. Detailed scaling analyses are performed on the distribution of coordination numbers. The fluctuation of the jamming density in finite-sized systems has important consequences on the finite-size effects of various quantities, resulting in a square relationship between disconnected and connected susceptibilities, a key signature of first-order transitions with quenched disorder. This study puts the jamming transition into the category of a broad class of transitions in disordered systems, including the first-order transition in the random-field Ising model, brittle yielding of amorphous materials, and melting of ultra-stable glasses.
Introduction.
Jamming in athermal particles is a paradigm of transitions between fluids and amorphous solids Makse et al. (2000); O’hern et al. (2003); Liu and Nagel (2010); van Hecke (2009); Pan et al. (2023a), with a deep connection to the glass transition in thermal systems Parisi and Zamponi (2010); Charbonneau et al. (2017); Parisi et al. (2020). Recent studies have revealed extremely rich features in the jamming phenomenon, but the nature of the jamming transition remains inconclusive.
(i) Mechanical marginality, related diverging length scales and power-law scalings. At the jamming transition density (volume fraction) $\varphi_{\rm J}$, the isostatic condition needs to be satisfied for the coordination number $Z$ (average number of contacts per particle): $Z=Z_{\rm iso}=2d$ in frictionless, infinite systems, where $d$ is the dimensionality Maxwell (1870); Alexander (1998). Isostaticity implies that at $\varphi_{\rm J}$ the system is marginally stable, inspiring the search for diverging length scales. According to the “cutting argument”, the isostaticity gives rise to a diverging isostatic length scale at jamming, $l^{*}\sim\Delta Z^{-1}\sim(\varphi-\varphi_{\rm J})^{-1/2}$, below which mechanical stability of the bulk system is affected by the cutting boundaries Wyart (2005); wya (2005); Silbert et al. (2005), where $\Delta Z=Z-Z_{\rm iso}$ is the excess coordination number. The effective medium theory gives a scattering length scale diverging at $\varphi_{\rm J}$, $l_{c}\sim\Delta Z^{-1/2}\sim(\varphi-\varphi_{\rm J})^{-1/4}$, below which continuum elasticity breaks down Wyart (2010); DeGiuli et al. (2014a); Lerner et al. (2014). Other related diverging length scales include the transverse wavelength, $\xi_{T}\sim(\varphi-\varphi_{\rm J})^{-\nu_{T}}$ with $\nu_{T}\approx 0.24$ Silbert et al. (2005), and the longitudinal wavelength, $\xi_{L}\sim(\varphi-\varphi_{\rm J})^{-\nu_{L}}$ with $\nu_{L}\approx 0.48$ Silbert et al. (2005).
At jamming, the marginal stability analysis provides relationships between the exponents $\theta$ appearing in the power-law distribution of weak inter-particle forces $P(f)\sim f^{\theta}$ and $\alpha$ in the distribution of small inter-particle gaps $P(h)\sim h^{-\alpha}$: $\theta=1/\alpha-2$ for extensive modes Wyart (2012) and $\theta=1-2\alpha$ for localized buckling modes Lerner et al. (2013); DeGiuli et al. (2014b). Above jamming, the marginal boundary between unstable and stable phases is defined by a scaling relation, $\Delta Z\sim(\varphi-\varphi_{\rm J})^{1/2}$ Wyart (2005); Müller and Wyart (2015). Other scalings have been established for over-jammed systems near $\varphi_{\rm J}$ Makse et al. (2000); O’hern et al. (2003); Goodrich et al. (2016); Liu and Nagel (2010); Baule et al. (2018); Pan et al. (2023a). For example, a relationship between the shear modulus $G$ and $\Delta Z$ can be derived by microscopic elastic theories, $G\sim\Delta Z\sim(\varphi-\varphi_{\rm J})^{1/2}$ Wyart (2005); Zaccone and Scossa-Romano (2011); Pan et al. (2023b).
(ii) Hyperuniformity and associated diverging length scales. Recent studies reveal the spatial distribution of the single-particle contact number, $Z_{i}$, is hyperuniform at jamming, implying a diverging hyperuniform correlation length, $\xi_{\rm H}\sim\Delta Z^{-\nu}$ Hexner et al. (2018, 2019). The value of exponent $\nu$ appears to depend on the way to extract the correlation length and dimensionality Hexner et al. (2018).
(iii) Gardner glass phase, landscape marginality and associated criticality. The marginality at jamming has been established by an independent approach within the framework of replica symmetry breaking Charbonneau et al. (2014); Parisi et al. (2020); Charbonneau et al. (2015a); Berthier et al. (2016, 2019); Li et al. (2021); Urbani et al. (2023). Unjammed hard sphere glasses undergo a Gardner transition where the free energy landscape becomes fractal and marginal, and the caging susceptibility diverges Berthier et al. (2016). The entire Gardner phase, including the jamming limit, is critical. In other words, the caging correlation length $\xi_{\rm G}\sim(\varphi_{\rm J}-\varphi)^{0}$ remains infinite near jamming. The mean-field replica theory predicts the values of exponents in the weak force and small gap distributions, $\alpha=0.41269$ and $\theta=0.42311$, coinciding with the relationship $\theta=1/\alpha-2$ given by the mechanical marginal stability analysis Charbonneau et al. (2014, 2015b); Wang et al. (2022). The theory provides an additional scaling relationship between the cage size $\Delta$ and the entropic pressure $p$, $\Delta\sim p^{-\kappa}$ with $\kappa=1.41574$. The exponents appear to be independent of the dimensionality $d$ for $d\geq d_{\rm u}$ Charbonneau et al. (2014, 2015b), where $d_{\rm u}=2$ is the conjectured upper critical dimension Goodrich et al. (2012).
(iv) Criticality in shear rheology. The criticality of the jamming transition is suggested by scaling analysis of the rheological data obtained in finite-rate shear simulations of flowing states near $\varphi_{\rm J}$ Olsson and Teitel (2007, 2011); Kawasaki et al. (2015). Combing the power-law divergence of the viscosity, $\eta=(\varphi_{\rm J}-\varphi)^{-\beta}$ and the vanishing of yield stress, $\sigma_{\rm Y}\sim(\varphi-\varphi_{\rm J})^{\Delta}$, a critical scaling function is proposed. A diverging rheological correlation length can be extracted from velocity correlations Olsson and Teitel (2007) or non-affine displacements Heussinger and Barrat (2009), $\xi_{\rm R}\sim(\varphi_{\rm J}-\varphi)^{-\nu}$, with $\nu=0.6-1.1$ Olsson and Teitel (2007, 2011); Heussinger and Barrat (2009); Vagberg et al. (2011). The criticality seems to be at odds with the hyperuniformity discussed in (ii) – in the thermodynamical limit, the fluctuations diverge in the former and vanish in the latter.
(v) Discontinuity in the coordination number. The coordination number $Z$, which is considered as an order parameter of the jamming transition, jumps discontinuously from $Z=0$ below jamming, to $Z\geq Z_{\rm iso}$ above, in quasi-static compression or decompression Makse et al. (2000); O’hern et al. (2003). Apparently, as the signature of a first-order transition, this discontinuity is inconsistent with the viewpoints of a continuous transition described above.
It is clear that various diverging length scales have been suggested throughout the literature. However, none of the lengths in (i-iv) can explain the finite-size scaling behavior of the jamming fraction $F_{\rm J}(\varphi,N)$ O’hern et al. (2003); Bertrand et al. (2016); Baity-Jesi et al. (2017); Jin and Yoshino (2021): the data can be reasonably collapsed by a master curve, $F_{\rm J}(\varphi,N)=\mathcal{F}[(\varphi-\varphi_{\rm J})N^{1/2}]$, valid for both compression and shear jamming in two (2D) and three (3D) dimensions. In this study, we show that this scaling can be fully explained by a first-order transition scenario of the jamming transition with quenched disorder. The form $(\varphi-\varphi_{\rm J})N^{1/2}$ originates from the disorder-induced fluctuation of the jamming density itself in finite-sized systems of $N$ particles, which follows the standard central limit theorem. Thus this finite-size scaling is independent of isostaticity, marginality, criticality and hyperuniformity.
Three important differences between previous approaches (i-v) and ours shall be denoted. First, in (i-iii), the jamming limit is approached from one side only, i.e., the over-jammed side ($\varphi>\varphi_{\rm J}$) in (i) and (ii), and the unjammed side ($\varphi<\varphi_{\rm J}$) in (iii). Here we consider a well-defined ensemble including both over-jammed and unjammed states, whose ratio is essential in the scaling analysis. Note that once an ensemble average is taken, the (v) discontinuity in $Z$ turns into a smooth function, and thus a finite-size analysis becomes essential to see the asymptotic behavior in the thermodynamic limit.
Second, in conventional compression quenching protocols, the generated ensemble depends on the initial conditions  Chaudhuri et al. (2010); Ozawa et al. (2012); Jin and Yoshino (2021) and the basins of attraction Bertrand et al. (2016); Ashwin et al. (2012); Martiniani et al. (2017). Here we instead consider an ensemble prepared by cyclic shear, where the states are sampled by well-controlled dynamics similar to that in thermal systems. Recently, the response of amorphous assemblies of particles to cyclic shear has attracted great interest, due to the presence of a nonequilibrium phase transition, called the reversible-irreversible (RI) transition Pine et al. (2005); Corte et al. (2008); Schreck et al. (2013); Nagasawa et al. (2019); Das et al. (2020). Interestingly, the jamming transition lies in the irreversible regime where particle trajectories are asymptotically diffusive Nagasawa et al. (2019); Das et al. (2020). In this study, we restrict our ensemble within the irreversible phase.
Third, in our ensemble, we carefully exclude partially crystallized and fragile states with $Z<Z_{\rm iso}$ that are sensitive to mechanical perturbations or protocol parameters, and regard them as unjammed states with $Z=0$. If such states are included, the discontinuity in $Z$ diminishes and the jamming transition looks continuous, similar to the results reported in Ref. Heussinger and Barrat (2009) obtained by uniform shear. Thus we expect the impactibility between the criticality viewpoint in (iv) and our first-order picture originating from finite-rate effects. For finite-time scales, the existence of transient states with $Z<Z_{\rm iso}$ can lead to a continuous jamming transition Olsson and Teitel (2007, 2011); Heussinger and Barrat (2009). However, our results suggest that, after a sufficiently long time, any configurations with $Z<Z_{\rm iso}$ would eventually relax to unjammed states with zero energy and inter-particle contacts.
An ensemble generated by cyclic athermal quasi-static shear.
We apply cyclic athermal quasistatic shear (CAQS) to standard models of soft, frictionless particles in 2D and 3D (see Models and Shear protocol in Methods). We present 2D data in the main text and 3D data in the Extended Data. For the 2D model, previous studies report a jamming density (J-point density) $\varphi_{\rm J}\approx 0.842-0.843$ O’Hern et al. (2002); Vagberg et al. (2011). The phase diagram of RI transitions near $\varphi_{\rm J}$ is plotted in Fig. 1a. The RI transition is characterized by the one-cycle displacement averaged over particles Nagasawa et al. (2019),
$\delta r(t)=\frac{1}{N}\sum_{i=1}^{N}\left|\boldsymbol{\mathrm{r}}_{i}(t+1)-\boldsymbol{\mathrm{r}}_{i}(t)\right|,$ (1)
and the mean-squared displacement (MSD),
$\langle\Delta r^{2}(t)\rangle=\frac{1}{N}\left\langle\sum_{i=1}^{N}\left|\boldsymbol{\mathrm{r}}_{i}(t)-\boldsymbol{\mathrm{r}}_{i}(0)\right|^{2}\right\rangle,$ (2)
where $t$ is the number of cycles playing a similar role as the time in thermal systems, $\boldsymbol{\mathrm{r}}_{i}(t)$ the position of particle $i$ at time $t$ and zero strain $\gamma=0$, and $\langle x\rangle$ the average over $N_{\rm s}$ samples.
The RI dynamics near $\varphi_{\rm J}$ are systematically studied in Ref. Nagasawa et al. (2019), according to which $\delta r(t)$ displays two-step relaxation behavior typically appearing in glassy systems. For $\tau_{\rm R}<t<\tau_{\rm L}$, $\delta r(t)$ develops a plateau at $\delta r_{\rm s}$. In the interested irreversible phase, $\delta r_{\rm s}>0$, $\tau_{\rm R}\sim 0$ and $\tau_{\rm L}>t_{\rm max}$ with $t_{\rm max}$ the maximum simulation time, suggesting that the system reaches a stationary state (see Fig. 1b). In addition, the MSD in the irreversible phase shows typical diffusive behavior $\langle\Delta r^{2}(t)\rangle\sim t$ (see Fig. 1c). In contrast, in the reversible phase, $\delta r_{\rm s}=0$, which means that the system is “absorbed” into an invariant state. In practice, one defines $\delta r_{\infty}=\delta r(t_{\rm max})$ and distinguish between the reversible and irreversible phases by comparing the value of $\delta r_{\infty}$ to a threshold $\delta r_{\rm th}$. In this study, we set $t_{\rm max}=4000$ and $\delta r_{\rm th}=0.1$, giving the boundary between irreversible (yellow) and reversible (blue) regions in Fig. 1a.
The above results imply that the configurational space is effectively explored by the dynamical trajectory of the system in the irreversible phase, encouraging us to consider a statistical ensemble generated by shear dynamics. In the main text below, we fix $\gamma_{\rm max}={0.7}$, and vary $\varphi$ systematically in the window of $[0.833,0.849]$. At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ independent configurations are collected to construct the ensemble. In the Extended Data, we present additional results obtained for $\gamma_{\rm max}={1}$. All configurations are typically isotropic since they are collected at $\gamma=0$.
Figure 1d shows the time evolution of the coordination number $Z(t)$ in the irreversible phase, obtained from a typical simulation near $\varphi_{\rm J}$. Rattlers (particles with fewer than $d+1$ contacts) are removed in the computation of $Z$. At first glance, one sees the coexistence of jammed ($Z\approx 2d=4$) and unjammed states ($Z=0$), similar to the coexistence of two ferromagnetic states (positive $m$ and negative $m$) in the time evolution of the magnetization $m(t)$ in an Ising model near a first-order phase transition Binder (1987a). However, a more careful examination reveals four distinct states, which we discuss below.
Figure 1: Cyclic athermal quasi-static shear simulations. (a) Reversible ($\delta r_{\infty}<0.1$, blue region) and irreversible ($\delta r_{\infty}>0.1$, yellow region) phases ($N={1000}$). The solid line represents $\varphi_{\rm J}^{\infty}={0.8432}$, and the dashed line represents $\gamma_{\rm max}=0.7$ used to generate the ensemble in the main text. (b) One-cycle displacement $\delta r(t)$ and (d) average coordination number $Z(t)$ for a typical sample. (c) The MSD data show diffusive behavior $\langle\Delta r^{2}(t)\rangle\sim t$ (line). Data in (b-c) are obtained at a fixed $\varphi=0.841$.
Four states: unjammed, jammed, partially crystallized and fragile.
Figure 2: Four states. (a) Semi-log plot of the probability distribution $p(Z)$. The four peaks of unjammed (U), partially crystallized (C), fragile (F) and jammed (J) states, are indicated. The inset shows a linear plot of $p(Z)$ without the unjammed peak. Fractions are plotted as a function of $\varphi$ in (b), and of $t$ in (c). In (c), each data point is obtained for a small window $[t-\delta t,t+\delta t]$ with $\delta t={250}$. (d) Peak coordination numbers $Z_{\rm J}^{*}$ of jammed states (upper branch, open symbols) and $Z_{\rm F}^{*}$ of fragile states (lower branch, filled symbols) as functions of $\varphi$, for $N=256$ (diamonds), 512 (triangles), 1000 (squares), and 2000 (circles). In (e-h), we show typical configurations of (e) partially crystallized, (f) fragile percolating in one direction, (g) fragile percolating in both directions, and (h) jammed states. Contact forces are represented by bonds, whose width is proportional to the magnitude of force. The red and blue disks are non-rattler and rattler particles respectively. We set $\varphi=0.841$ for (a) and (c), $N=256$ for (b) and (c), and $N=1000$ for (e-h).
In Fig. 2a, we plot the probability distribution $p(Z)$ of the states in the considered ensemble, at a fixed $\varphi=0.841$ in the irreversible phase. The following four states, represented by peaks in $p(Z)$, can be identified (see Figs. 2e-h).
(i) Unjammed states. The left-most peak is a delta-function $p_{\rm U}(Z)\sim\delta(Z)$, corresponding to unjammed states. All unjammed states have strictly zero contacts, $Z=0$, once rattlers are removed.
(ii) Jammed states. The right-most peak $p_{\rm J}(Z)$ at $Z\geq 4$ corresponds to jammed states. Their average coordination numbers satisfy minimally the isostatic condition, $Z\geq 2d=4$.
(iii) Partially crystallized states. The delta-peak at $p_{\rm C}(Z)\sim\delta(Z-24/7)$ represents the states with a single unit cell of the hexagonal crystal (see Fig. 2e). Here $24/7\approx 3.4286$ is the average number of contacts of the seven particles forming the unit cell. Occasionally, states with two or three unit crystal cells can be also found but they are rare.
(iv) Fragile states. We define the states in the broad peak $43<Z<$ 4, excluding the crystalline peak $p_{\rm C}(Z)$ as the fragile states.
The fractions, $F_{\rm U}$, $F_{\rm J}$, $F_{\rm C}$ and $F_{\rm F}$, of the above four states (i-iv), are obtained by integrating corresponding peaks in $p(Z)$. In Fig. 2b, the fractions are plotted as functions of $\varphi$, showing that $F_{\rm J}(\varphi)$ increases from zero to one across $\varphi_{\rm J}$. This behavior is quantitatively similar to $F_{\rm J}(\varphi)$ obtained in previous rapid quench simulations, where finite-size analyses have been carried out to precisely determine the asymptotic jamming transition density $\varphi_{\rm J}^{\infty}$ in the thermodynamic limit $N\rightarrow\infty$ O’hern et al. (2003); Vagberg et al. (2011). We will perform such finite-size analyses later, after discussing the nature of fragile states.
In Fig. 2c, we report various fractions as functions of $t$ at $\varphi={0.841}$. The fractions are independent of $t$, confirming that the system is stationary. We further divide $F_{\rm C}(t)$ into two parts, $F_{\rm C}(t)=F_{\rm C}^{1}(t)+F_{\rm C}^{>}(t)$, where $F_{\rm C}^{1}(t)$ and $F_{\rm C}^{>}(t)$ are respectively the fractions of partially crystallized states with one and more than one crystal unit cells. Both $F_{\rm C}^{1}(t)$ and $F_{\rm C}^{>}(t)$ are independent of $t$, indicating that the growth of seed crystals is not observed. Because $F_{\rm C}$ is generally several orders of magnitude lower than the fractions of other types, partially crystallized states will be ignored in the following analyses.
In Fig. 2d, we plot the maximum coordination numbers, $Z_{\rm J}^{*}$ and $Z_{\rm U}^{*}$, of jammed and fragile peaks. Around $\varphi_{\rm J}$, there is a small gap $\delta Z_{\rm gap}\approx 0.2$ between $Z_{\rm J}^{*}$ (upper branch) and $Z_{\rm U}^{*}$ (lower branch). The results are very similar to those obtained by quasistatic uniform shear in Ref. Heussinger and Barrat (2009) (see Fig. 4 therein). In Ref. Heussinger and Barrat (2009), it is suggested that $\delta Z_{\rm gap}$ vanishes in the thermodynamic limit, and thus the jamming transition is continuous. However, as we show below, the fragile states are generated due to incomplete energy minimization and are mechanically unstable. Once such fragile states are excluded, the lower branch only contains unjammed states with $Z_{\rm U}^{*}=0$, which is separated by a large gap $\delta Z_{\rm gap}\approx 4$ from the upper branch $Z_{\rm J}^{*}\geq 4$.
Instability of fragile states.
Previously, the fragile states were obtained by uniform shear at non-zero strains $\gamma>0$ below $\varphi_{\rm J}$ in experiments Bi et al. (2011) and simulations Kumar and Luding (2016). It was proposed that fragile and jammed states differ in the percolation of the strong force network: in fragile states the percolation occurs anisotropically in one direction only, while in jammed states it occurs isotropically in both directions. In this study, all states are collected during cyclic shear at $\gamma=0$, without systematically introducing anisotropy. Indeed, the fragile states with $43<Z<$ 4 can have percolated force networks of non-rattlers in one or two directions (see Fig. 2f and g). Thus in our case, anisotropy cannot effectively distinguish fragile from jammed states.
We demonstrate that the essential difference between fragile and jammed states is their mechanical stability. In fragile states, the potential energy per particle is negligibly low (see Stopping criterion for energy minimization in Methods), $\overline{e}_{i}<e_{\rm th}=10^{-20}$ (see Fig. 3a), but the fraction of non-rattler particles is non-zero. These non-rattlers experience forces, which can form a transient, percolated network. Such networks are highly heterogeneous (see Fig. 2f and g), compared to those in jammed states (see Fig. 2h). More importantly, the force networks in fragile states are unstable, revealed by the non-zero net force per particle, $\overline{f}_{i}>f_{\rm th}=10^{-14}$ (see Fig. 3b). Thus fragile states are not strictly equilibrated; they turn into unjammed states by sufficiently long relaxation (accurate energy minimization) or mechanical perturbations.
To demonstrate the instability of fragile states, two tests are carried out. First, we perform a compression-decompression perturbation, $\varphi\rightarrow\varphi+\delta\varphi\rightarrow\varphi$, where $\delta\varphi=10^{-8}$, and the energy is minimized after each step. All fragile states ($43<Z<$ 4) become unjammed ($Z=0$) after this perturbation, while jammed states remain. In Fig. 3c, we plot the distribution $p^{\prime}(Z)$ after the compression-decompression perturbation. Independently, without any perturbation we simply regard all fragile states ($43<Z<$ 4) as unjammed ($Z=0$) and recalculate the distribution $\tilde{p}(Z)$. Figure 3c shows that the two distributions $p^{\prime}(Z)$ and $\tilde{p}(Z)$ perfectly coincide.
In the second test, we repeat CAQS simulations by systematically varying the threshold $f_{\rm th}$ in the criterion of the energy minimization algorithm. The fraction $F_{\rm F}$ of fragile states grows and $F_{\rm U}$ of unjammed states decays with increasing $f_{\rm th}$ (Fig. 3d), suggesting that many unjammed states become fragile under a looser force-balance condition $\overline{f}_{i}<f_{\rm th}$ with a larger $f_{\rm th}$. In contrast, the constant $F_{\rm J}$ shows that the definition of jammed states is insensitive to the algorithm parameter. Note that in granular experiments Bi et al. (2011); Zhao et al. (2019), the friction may play the role of $f_{\rm th}$, i.e., the net inter-particle force could be balanced by the frictional force between particles and the supporting plate. According to this assumption, the probability of observing fragile states in experiments would depend on the particle-plate friction that can be changed by replacing the materials. It provides a protocol to examine our scenario of fragile states in future experimental studies.
Because fragile states are unstable, from now on we count them as unjammed states. More specifically, we replace the original distribution $p(Z)$ with the modified distribution $\tilde{p}(Z)$. For simplicity, we omit the tilde and denote $\tilde{p}(Z)$ by $p(Z)$ below.
Figure 3: Stability tests. Average single-particle (a) potential energy $\overline{e}_{i}$ and (b) net force $\overline{f}_{i}$ as functions of $Z$. Lines in (a) and (b) are guides to the eye. (c) Comparison between $p^{\prime}(Z)$ obtained after the compression-decompression perturbation and $\tilde{p}(Z)$ with fragile states counted as unjammed states, for $N=1000$ at $\varphi=0.844$. (d) Fractions of four states as functions of $f_{\rm th}$. Data are obtained for $N=256$ systems, except for (c).
Figure 4: Distribution of the coordination number. (a) Probability distribution $p(Z)$ at a few different $\varphi$ for $N=1000$. (b) $p(Z)$ at $\varphi=0.843$ for a few different $N$. For better visualization, we do not show the unjammed delta-peak at $Z=0$. (c) The fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8432$. (d) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid line represents fitting to Eq. (4), with two fitting parameters $u=0.041$ and $\sigma_{\varphi}=0.043$. The average coordination number $\langle Z\rangle$ is plotted as a function of $\varphi$ in (e) and of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$ in (f). The solid line in (f) represents $\langle Z\rangle=F_{\rm J}Z_{\rm J}^{*}$ using $u=0.041$, $\sigma_{\varphi}=0.043$ and $Z_{\rm J}^{*}=4.1$. Symbols in (b-f) have the same meanings.
Scaling analysis of $p(Z)$ near the jamming transition.
The new distribution $p(Z)$ contains two peaks of unjammed and jammed states, whose fractions satisfy $F_{\rm U}(\varphi)+F_{\rm J}(\varphi)=1$ at any $\varphi$. In Figs 4a and b, we plot $p(Z)$ for a few different $\varphi$ and $N$, showing that $p(Z)$ is always double-peaked across the jamming transition, which is typical behavior of a first-order rather than a second-order transition.
To analyze the scaling behavior of $p(Z)$, we consider a general first-order form of first-order phase transitions Binder (1987a, b),
$p(Z)=(1-F_{\rm J})\delta(Z)+F_{\rm J}\,p_{\rm J}\left[(Z-Z_{\rm J}^{*})N^{\eta}\right],$ (3)
where $\delta(Z)$ and $p_{\rm J}(Z)$ correspond to the unjammed and jammed peaks respectively. The asymptotic jamming density $\varphi_{\rm J}^{\rm\infty}=0.8432$ for $N\rightarrow\infty$ is determined by the intersection of $F_{\rm J}(\varphi)$ curves with different $N$ (see Fig. 4c), close to the J-point density $\varphi_{\rm J}\approx 0.842-0.843$ reported in previous studies O’hern et al. (2003); Vagberg et al. (2011). The difference between $\varphi_{\rm J}^{\rm\infty}=0.8432$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}=0.8435$ by $\gamma_{\rm max}=1$ (see Extended Data Fig. E1) is too small to conclude any systematic dependence on $\gamma_{\rm max}$. In Ref. Das et al. (2020), an unjamming pocket is reported in the phase diagram for $\gamma_{\rm max}<0.17$, while for larger $\gamma_{\rm max}$, $\varphi_{\rm J}$ is independent of $\gamma_{\rm max}$, consistent with our observations. These results also suggest that the jamming density obtained by cyclic shear is the lowest jamming density on the J-line of all possible jammed states Parisi and Zamponi (2010); Chaudhuri et al. (2010); Babu et al. (2021).
As shown in Fig. 2d, $Z_{\rm J}^{*}$ is independent of $N$ near $\varphi_{\rm J}$, and weakly depends on $\varphi$. In the following scaling analysis, $Z_{\rm J}^{*}$ is approximated by a constant value $Z_{\rm J}^{*}\approx 4.1$ at $\varphi_{\rm J}^{\rm\infty}=0.8432$. To keep the expressions concise, we ignore the corrections to the scaling functions from the $\varphi$-dependence of $Z_{\rm J}^{*}$.
We assume the fraction of jammed states $F_{\rm J}$ having the following scaling form, $F_{\rm J}(\varphi,N)=\mathcal{F}\left[(\varphi-\varphi_{\rm J}^{\infty})N^{\lambda}\right]$. The value of the exponent $\lambda$ is important for understanding the nature of the transition. If the system were thermal, the fraction would be determined by the Boltzmann distribution, $F_{\rm J}\sim\exp\left(\frac{N\delta f}{k_{\rm B}T}\right)$, where $\delta f$ is the single-particle free energy difference between two phases Binder (1987b, a). Because the free energy is non-singular around a first-order phase transition, it can be expanded, giving $\delta f\sim(T-T_{\rm c})$ to the lowest order. Thus, if the jamming transition were a standard first-order phase transition, then $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}\left[(\varphi-\varphi_{\rm J}^{\infty})N\right]$, i.e., $\lambda=1$ (note that $\varphi$ in the granular system is analogous to the temperature $T$ in thermal systems). However, our numerical data can not be collapsed using $\lambda=1$; in contrast, they can be perfectly collapsed using $\lambda=1/2$ (see Fig. 4d).
To understand the origin of $\lambda=1/2$ scaling, recall that the athermal jamming transition is not driven by the free energy difference between the two phases. Below we explain the scaling by the scenario of a first-order transition with quenched disorder. For a finite $N$, due to the presence of disorder in the packing structure, the jamming density $\varphi_{\rm J}^{N}$ should fluctuate around the asymptotic value $\varphi_{\rm J}^{\infty}$. Let us assume a simple Gaussian form of the distribution, $\rho(\varphi_{\rm J}^{N})\sim\exp\left[-\frac{(\delta\hat{\varphi}_{\rm J}+u)^{2}}{2\sigma_{\varphi}^{2}}\right],$ where $\delta\hat{\varphi}_{\rm J}=(\varphi_{\rm J}^{N}-\varphi_{\rm J}^{\infty})N^{1/2}$ follows the standard central limit theorem. Note that $\rho(\varphi_{\rm J}^{N})$ has been explicitly measured in the compression protocol O’Hern et al. (2002); O’hern et al. (2003): the width of $\rho(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-0.55}$ in both 2D and 3D, supporting our assumption. The above assumption also predicts, $\varphi_{\rm J}^{\infty}-\varphi_{\rm J}^{N}\sim N^{-1/2}$, independent of dimensionality. The jammed states are defined by $\varphi>\varphi_{\rm J}^{N}$, and thus
[…]
where $\delta\hat{\varphi}=(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. Equation (4) agrees well with the data in Fig. 4d.
Figure 5: Susceptibilities. We plot (a) the disconnected susceptibility $\chi_{\rm dis}$ rescaled by $N$, (b) the connected susceptibility $\chi_{\rm con}$ rescaled by $N^{1/2}$, and (c) the ratio between $\chi_{\rm dis}$ and $\chi_{\rm con}^{2}$, as functions of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid black lines in (a-c) are Eqs. (5), (6) and (7) respectively. The dashed red line in (c) is the constant term in Eq. (7), i.e., $\chi_{\rm dis}/\chi_{\rm con}^{2}=\pi\sigma_{\varphi}^{2}/2$. To draw these lines, we use $u=0.036$ and $\sigma_{\varphi}=0.043$ obtained from the fitting in Fig. 4d, and $Z_{\rm J}^{*}=4.1$ determined at $\varphi_{\rm J}^{\infty}=0.8432$ in Fig. 2d. No fitting is performed here. Symbols in (a-c) have the same meanings as indicated in the legend of (a). In (d) and (e), the susceptibilities of jammed states $\chi_{\rm J}$ are plotted as functions of $N$ for a few different $\varphi$, and are fitted to Eq. (8). The open squares in (d) are $\chi_{\rm J}$ data obtained in a small pressure window $0.00110.0009<P<$ 0.0011, corresponding to $\varphi\approx 0.843$. The fitting parameter $\mu$ is plotted in (f) as a function of $\varphi$, where the error bar represents the fitting error.
[…]
$\frac{\chi_{\rm dis}(\varphi,N)}{N}\approx(Z_{\rm J}^{*})^{2}[1-F_{\rm J}(\varphi,N)]F_{\rm J}(\varphi,N),$ (5)
where $F_{\rm J}(\varphi,N)$ is given by Eq. (4). Equation (5) suggests a scaling form $\chi_{\rm dis}(\varphi,N)/N=\mathcal{X}_{\rm dis}\left[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}\right]$, in agreement with the data in Fig. 5a.
The alternative connected susceptibility is defined as $\chi_{\rm con}=d\langle Z\rangle/d\varphi$. From Eq. (3), one finds that $\langle Z\rangle\approx F_{\rm J}Z_{\rm J}^{*}$ and thus $\langle Z\rangle$ has the same scaling form as $F_{\rm J}$, i.e., $\langle Z\rangle(\varphi,N)=\mathcal{Z}[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}]$, consistent with the data in Fig. 4e and f. Using $\chi_{\rm con}=Z_{\rm J}^{*}\frac{dF_{\rm J}}{d\varphi}$ and Eq. (4), we derive,
$\frac{\chi_{\rm con}(\varphi,N)}{N^{1/2}}\approx\frac{Z_{\rm J}^{*}}{\sigma_{\varphi}\sqrt{2\pi}}\exp\left[-\left(\frac{\delta\hat{\varphi}+u}{\sqrt{2}\sigma_{\varphi}}\right)^{2}\right],$ (6)
which is verified by the simulation data in Fig. 5b.
Now we can look at the relationship between $\chi_{\rm dis}$ and $\chi_{\rm con}$. Expanding Eqs. (5) and  (6) around the maxima, where $x\equiv\frac{\delta\hat{\varphi}+u}{\sqrt{2}\sigma_{\varphi}}=0$, we obtain, up to the quadratic order,
$\frac{\chi_{\rm dis}}{\chi_{\rm con}^{2}}\approx\frac{\pi\sigma^{2}_{\varphi}}{2}\left[1+\left(2-\frac{4}{\pi}\right)\left(\frac{\delta\hat{\varphi}+u}{\sqrt{2}\sigma_{\varphi}}\right)^{2}\right].$ (7)
To the lowest order, Eq. (7) gives a scaling, $\chi_{\rm dis}\sim\chi_{\rm con}^{2},$ which is a key signature of the presence of quenched disorder. This square relationship appears generally at first-order transitions in disordered systems, such as the first-order phase transition in the RFIM Nattermann (1998), the brittle yielding of amorphous solids Ozawa et al. (2018) and the melting of ultrastable hard-sphere glasses Zhang et al. (2022). Comparing to the simulation data, Eq. (7) works well around the extreme point (see Fig. 5c).
It seems that the non-Gaussian effect in the distribution $\rho(\varphi_{\rm J}^{N})$, which is neglected in the present analysis, is amplified in the data of $\chi_{\rm dis}/\chi_{\rm con}^{2}$, resulting in slight asymmetry. Extended Data Figs. E1-E4 show that the scaling functions Eqs. (4)-(7) work in 3D, and for a different max strain $\gamma_{\max}=1$ in 2D.
Finally let us discuss the finite-size effects of the jammed peak $p_{\rm J}(Z)$ in Eq. (3). For standard first-order phase transitions, $\eta=1/2$ Binder (1987b, a). Thus in that scenario the susceptibility of the jamming peak would scale as, $\chi_{\rm J}\equiv N\sigma_{\rm J}^{2}=N\langle(Z-Z_{\rm J}^{*})^{2}\rangle_{\rm J}\sim N^{0}$, where $\sigma_{\rm J}^{2}$ is the variance of $p_{\rm J}(Z)$ and $\langle\ldots\rangle_{\rm J}$ represents the average restricted to the jammed states only. However, our $\chi_{\rm J}$ data disagree with this scaling (see Figs.  5d-f). At different $\varphi$, $\chi_{\rm J}$ can be fitted to a pow-law form,
$\chi_{\rm J}\sim N^{-\mu/d},$ (8)
where $\mu=d(2\eta-1)$ is a fitting parameter. At large or small $\varphi$ away from $\varphi_{\rm J}^{\rm\infty}$ (see Fig.  5f), the exponent is close to zero ($\mu\approx-0.2$), suggesting that the local coordination numbers $Z_{i}$ are uniformly distributed. The deviation from the uncorrelated behavior ($\mu=0$) is significant around $\varphi_{\rm J}^{\rm\infty}$, where $\mu\approx-0.8$ reaches the minimum.
The result of $\mu$ shall be interpreted with care. In general, (i) $\mu=0$ corresponds to uniformity of $Z_{i}$, (ii) $\mu>0$ to hyperuniformity with a vanishing $\chi_{\rm J}$ in the thermodynamic limit, and (iii) $\mu<0$ to hyperfluctuations with a diverging $\chi_{\rm J}$ in the thermodynamic limit that typically appears at the critical point in a second-order phase transition. However, the negative $\mu$ in Fig. 5f is not due to the criticality of the jamming transition. Here it is essential to consider the volume fluctuations. The negative $\mu$ is obtained from a $\varphi$-controlled setup in our CAQS simulations. If we instead select configurations around a constant pressure $P$, then the fluctuations become significantly smaller, and $\mu\approx 0$ (Fig. 5d). This observation is consistent with a previous study Ikeda (2023), suggesting that the large fluctuation of $Z$ near $\varphi_{\rm J}^{\rm\infty}$ in the $\varphi$-controlled protocol might be originated from the fluctuation of $\varphi_{\rm J}^{N}$. In fact, careful measurement gives $\mu\simeq 1$ in a $P$-controlled protocol at jamming, confirming hyperuniformity Hexner et al. (2018, 2019). We emphasize that the finite-size effects of the jammed peak $p_{\rm J}(Z)$ contribute negligibly to the scaling of $F(\varphi,N)$, $\chi_{\rm dis}(\varphi,N)$ and $\chi_{\rm con}(\varphi,N)$. In other words, the first-order nature of the jamming transition is independent of how $Z_{i}$ is spatially distributed in jammed packings.
Conclusion.
In this study, we understand the nature of jamming transition through an ensemble approach analogous to the statistical mechanics in equilibrium systems Baule et al. (2018); Bi et al. (2015). Within such a framework, the order of the jamming transition is independent of the complex properties of jammed packings, including isostaticity, marginality and hyperuniformity (see (i-iii) in Introduction). As a first-order jamming transition, the finite-size scaling exponents measured in this study are naturally universal in 2D and 3D, without the need to introduce the upper critical dimension $d_{\rm u}=2$ Parisi et al. (2018), a concept relevant to continuous transitions only. It is of particular interest to reconcile the first-order transition established here under quasi-static shearing and the second-order transition observed in previous finite-rate rheology (see (iv) in Introduction). The jamming transition appears to be a rare example with the simultaneous existence of first-order and second-order transitions between liquid and solid states. Conventionally, the two kinds of transitions only coexist in gas-liquid systems, but not in liquid-crystalline solid systems.
Methods.
Models. We study models of frictionless, bidisperse particles in two and three dimensions. The number ratio between large and small particles is 1:1, and the diameter ratio is 1.4:1. The potential energy between two particles is:
$U(r_{ij})=\frac{\epsilon}{2}\left(1-\frac{r_{ij}}{\sigma_{ij}}\right)^{2}\Theta\left(1-\frac{r_{ij}}{\sigma_{ij}}\right),$ (9)
where $\epsilon=1$ is the energy unit, $r_{ij}$ the distance between particles $i$ and $j$, $\sigma_{ij}=\frac{\sigma_{i}+\sigma_{j}}{2}$ the mean diameter, and $\Theta(x)$ the Heaviside step function. We set unit particle mass, and the diameter of small particles as unit length.
Shear protocol. Particles are randomly distributed at an initial volume fraction $\varphi_{0}=0.02$. The system is then rapidly quench compressed to the target $\varphi$. The CAQS is performed under the Lees-Edwards boundary conditions, with a fixed $\varphi$. We use a strain step $\delta\gamma=0.1$ to generate the phase diagram in Fig. 1a, and $\delta\gamma=0.01$ for other results. At each step, particle positions are shifted according to $x_{i}\rightarrow x_{i}+\delta\gamma y_{i}$, and then the system’s energy is minimized using the FIRE algorithm Bitzek et al. (2006). During a cycle, the strain is varied as $\gamma=0\rightarrow\gamma_{\rm max}\rightarrow 0$. We set $\gamma_{\rm max}=0.7$ in the main text and $\gamma_{\rm max}=1$ in the Extended Data for the 2D model, and $\gamma_{\rm max}=0.5$ for the 3D model (Extended Data). The number of cycles is represented by $t$ with unit oscillation period ($t=1$). In 2D, the maximum number of cycles is $t_{\rm max}=250$ for $0.839\leq\varphi\leq{0.849}$ and $t_{\rm max}=4000$ for $0.839{0.833}\leq\varphi<$ 0.839, while in 3D, $t_{\rm max}=144$ for $2000$, and $t_{\rm max}=50$ for $N=512$ and $1000$. We generate $N_{\rm s}=4000-12000$ independent samples at each $\varphi$. At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ configurations are collected to compute statistical quantities.
Stopping criterion for energy minimization. We terminate the energy minimization when the potential energy per particle $\overline{e}_{i}=\frac{1}{N}\sum_{i=1}^{N}e_{i}$ falls below a threshold $e_{\rm th}$ or the average single-particle net force $\overline{f}_{i}=\frac{1}{N}\sum_{i=1}^{N}f_{i}$ falls below $f_{\rm th}$, whichever is satisfied earlier. We set $e_{\rm th}=10^{-20}$ and $f_{\rm th}=10^{-14}$ unless otherwise specified.
Acknowledgements.
We warmly thank Hajime Yoshino for inspiring discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 11974361 and 12047503), from Chinese Academy of Sciences (Grants ZDBS-LY-7017 and KGFZD-145-22-13), and from Wenzhou Institute (Grant WIUCASQD2023009). In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.

### v2

@@
 The jamming transition seems to display mixed aspects of a first-order transition, evidenced by a discontinuity in the coordination number, and a second-order transition, indicated by power-law scalings and diverging lengths.
-Here we demonstrate that jamming is a first-order transition with quenched disorder in cyclically sheared systems with quasistatic deformations.
-Particle models are simulated in two and three dimensions, which undergo reversible-irreversible transitions under cyclic shear.
-The ensemble of configurations, with partially crystallized and fragile states excluded, is generated in the irreversible phase, where the system is stationary and particles are diffusive.
-Detailed scaling analyses are performed on the distribution of coordination numbers.
-The fluctuation of the jamming density in finite-sized systems has important consequences on the finite-size effects of various quantities, resulting in a square relationship between disconnected and connected susceptibilities, a key signature of first-order transitions with quenched disorder.
-This study puts the jamming transition into the category of a broad class of transitions in disordered systems, including the first-order transition in the random-field Ising model, brittle yielding of amorphous materials, and melting of ultra-stable glasses.
+Here we demonstrate that jamming is a first-order transition with quenched disorder in cyclically sheared systems with quasistatic deformations, in two and three dimensions.
+Based on scaling analyses, we show that fluctuations of the jamming density in finite-sized systems have important consequences on the finite-size effects of various quantities, resulting in a square relationship between disconnected and connected susceptibilities, a key signature of the first-order transition with quenched disorder.
+This study puts the jamming transition into the category of a broad class of transitions in disordered systems where sample-to-sample fluctuations dominate over thermal fluctuations, suggesting that the nature and behavior of the jamming transition can be understood within the developed theoretical framework of the athermally driven random-field Ising model.
 Introduction.
@@
 (ii) Hyperuniformity and associated diverging length scales.
-Recent studies reveal the spatial distribution of the single-particle contact number, $Z_{i}$, is hyperuniform at jamming, implying a diverging hyperuniform correlation length, $\xi_{\rm H}\sim\Delta Z^{-\nu}$ Hexner et al.
-(2018, 2019).
-The value of exponent $\nu$ appears to depend on the way to extract the correlation length and dimensionality Hexner et al.
-(2018).
+Recent studies reveal that the spatial distribution of the single-particle contact number, $Z_{i}$, is hyperuniform at jamming.
+The contact hyperuniformity is established by two power-law scalings measured in simulated packings Hexner et al.
+(2018, 2019): the scaling of the fluctuations of the average contact number in a hyper-cube of volume $\ell^{d}$, $\sigma_{Z}(\ell)\sim\ell^{-\mu}$ with $\mu=1$, and the small wave-vector scaling of the contact number structure factor, $S_{Z}(q)\sim q^{\alpha_{Z}}$ with $\alpha_{Z}\approx 1.53$.
+The crossover from the hyperuniform regime ($\ell\ll\xi_{\rm H}$) to the uniform regime ($\ell\gg\xi_{\rm H}$) defines a hyperuniform correlation length $\xi_{\rm H}$, which diverges at the jamming transition, $\xi_{\rm H}\sim\Delta Z^{-\nu_{\rm H}}$.
+The value of the exponent $\nu_{\rm H}$ appears to depend on how to extract the correlation length and the dimensionality.
 (iii) Gardner glass phase, landscape marginality and associated criticality.
@@
 The entire Gardner phase, including the jamming limit, is critical.
-In other words, the caging correlation length $\xi_{\rm G}\sim(\varphi_{\rm J}-\varphi)^{0}$ remains infinite near jamming.
+In other words, the caging correlation length, $\xi_{\rm G}\sim\infty$, remains infinite near jamming, when the transition is approached ($\varphi\rightarrow\varphi_{\rm J}$) from below ($\varphi<\varphi_{\rm J}$).
 The mean-field replica theory predicts the values of exponents in the weak force and small gap distributions, $\alpha=0.41269$ and $\theta=0.42311$, coinciding with the relationship $\theta=1/\alpha-2$ given by the mechanical marginal stability analysis Charbonneau et al.
@@
 The exponents appear to be independent of the dimensionality $d$ for $d\geq d_{\rm u}$ Charbonneau et al.
-(2014, 2015b), where $d_{\rm u}=2$ is the conjectured upper critical dimension Goodrich et al.
+(2014, 2015b); Parisi et al.
+(2018), where $d_{\rm u}=2$ is the conjectured upper critical dimension of the jamming transition Goodrich et al.
 (2012).
 (iv) Criticality in shear rheology.
-The criticality of the jamming transition is suggested by scaling analysis of the rheological data obtained in finite-rate shear simulations of flowing states near $\varphi_{\rm J}$ Olsson and Teitel (2007, 2011); Kawasaki et al.
+The criticality of the jamming transition is suggested by scaling analyses of the rheological data obtained in finite-rate shear simulations of flowing states near $\varphi_{\rm J}$ Olsson and Teitel (2007, 2011); Kawasaki et al.
 (2015).
 Combing the power-law divergence of the viscosity, $\eta=(\varphi_{\rm J}-\varphi)^{-\beta}$ and the vanishing of yield stress, $\sigma_{\rm Y}\sim(\varphi-\varphi_{\rm J})^{\Delta}$, a critical scaling function is proposed.
-A diverging rheological correlation length can be extracted from velocity correlations Olsson and Teitel (2007) or non-affine displacements Heussinger and Barrat (2009), $\xi_{\rm R}\sim(\varphi_{\rm J}-\varphi)^{-\nu}$, with $\nu=0.6-1.1$ Olsson and Teitel (2007, 2011); Heussinger and Barrat (2009); Vagberg et al.
-(2011).
+A diverging rheological correlation length can be extracted from velocity correlations Olsson and Teitel (2007) or non-affine displacements Heussinger and Barrat (2009), $\xi_{\rm R}\sim(\varphi_{\rm J}-\varphi)^{-\nu}$, with $\nu\approx 1$  Olsson and Teitel (2007); Vagberg et al.
+(2011); Olsson and Teitel (2020).
 The criticality seems to be at odds with the hyperuniformity discussed in (ii) – in the thermodynamical limit, the fluctuations diverge in the former and vanish in the latter.
 (v) Discontinuity in the coordination number.
-The coordination number $Z$, which is considered as an order parameter of the jamming transition, jumps discontinuously from $Z=0$ below jamming, to $Z\geq Z_{\rm iso}$ above, in quasi-static compression or decompression Makse et al.
+The coordination number $Z$, which is considered as an order parameter of the jamming transition Liu and Nagel (2010), jumps discontinuously from $Z=0$ below jamming, to $Z\geq Z_{\rm iso}$ above, under quasi-static compression/decompression Makse et al.
+(2000); O’hern et al.
+(2003) or shear Vinutha et al.
+(2020).
+In unjammed states, $Z=0$ because particles can push each other apart, leaving no overlapping between them.
+In jammed states, the contact network spans the entire system and the minimum condition to have such networks is $Z=Z_{\rm iso}$.
+Thus the discontinuous jump in $Z$ at jamming is essentially related to isostaticity.
+Apparently, as the signature of a first-order transition, this discontinuity is inconsistent with the viewpoint of a continuous transition described above.
+Note that not all physical quantities exhibit a discontinuity at jamming: for example, the pressure $P$ and the shear modulus $G$ vanish continuously when the jamming transition is approached from above Makse et al.
 (2000); O’hern et al.
 (2003).
-Apparently, as the signature of a first-order transition, this discontinuity is inconsistent with the viewpoints of a continuous transition described above.
+On the other hand, other quantities, such as the bulk modulus $B$ and the fraction of non-rattlers $f_{\rm NR}$, jump abruptly at jamming, similar to $Z$.
 It is clear that various diverging length scales have been suggested throughout the literature.
@@
 (2016); Baity-Jesi et al.
-(2017); Jin and Yoshino (2021): the data can be reasonably collapsed by a master curve, $F_{\rm J}(\varphi,N)=\mathcal{F}[(\varphi-\varphi_{\rm J})N^{1/2}]$, valid for both compression and shear jamming in two (2D) and three (3D) dimensions.
+(2017); Jin and Yoshino (2021): the data can be reasonably collapsed by a scaling form, $F_{\rm J}(\varphi,N)=\mathcal{F}[(\varphi-\varphi_{\rm J})N^{1/2}]$, valid for both compression and shear jamming in two (2D) and three (3D) dimensions.
 In this study, we show that this scaling can be fully explained by a first-order transition scenario of the jamming transition with quenched disorder.
@@
 Note that once an ensemble average is taken, the (v) discontinuity in $Z$ turns into a smooth function, and thus a finite-size analysis becomes essential to see the asymptotic behavior in the thermodynamic limit.
-Second, in conventional compression quenching protocols, the generated ensemble depends on the initial conditions  Chaudhuri et al.
+Second, in conventional compression protocols, the generated ensemble depends on the initial conditions  Chaudhuri et al.
 (2010); Ozawa et al.
@@
 (2017).
-Here we instead consider an ensemble prepared by cyclic shear, where the states are sampled by well-controlled dynamics similar to that in thermal systems.
+Here we instead consider an ensemble prepared by cyclic shear, where the states are sampled by well-controlled dynamics similar to those in thermal systems.
 Recently, the response of amorphous assemblies of particles to cyclic shear has attracted great interest, due to the presence of a nonequilibrium phase transition, called the reversible-irreversible (RI) transition Pine et al.
@@
 In this study, we restrict our ensemble within the irreversible phase.
+The jamming density of this ensemble is equivalent to the minimum jamming density, or the jamming-point (J-point) density that is obtained by rapid quench O’hern et al.
+(2003).
 Third, in our ensemble, we carefully exclude partially crystallized and fragile states with $Z<Z_{\rm iso}$ that are sensitive to mechanical perturbations or protocol parameters, and regard them as unjammed states with $Z=0$.
@@
 However, our results suggest that, after a sufficiently long time, any configurations with $Z<Z_{\rm iso}$ would eventually relax to unjammed states with zero energy and inter-particle contacts.
+Note that the timescale for the system to attain force balance diverges at the shear jamming transition Vinutha et al.
+(2020), which means that the real quasistatic limit would correspond to extremely small shear rates in large systems.
+Figure 1: Cyclic athermal quasi-static shear simulations.
+(a) Reversible ($\delta r_{\infty}<0.1$, blue region) and irreversible ($\delta r_{\infty}>0.1$, yellow region) phases ($N={1000}$).
+The solid line represents $\varphi_{\rm J}^{\infty}={0.8432}$, and the dashed line represents $\gamma_{\rm max}=0.7$ used to generate the ensemble in the main text.
+(b) One-cycle displacement $\delta r(t)$ and (d) coordination number $Z(t)$ for a typical sample, at $\varphi=0.841$.
+(c) Several typical particle trajectories during 100 cycles at $\varphi=0.8425$.
+Positions are recorded at $\gamma=0$ only, and the system size is $43.8\times 43.8$ with periodic boundary conditions.
+(inset) The MSD data show diffusive behavior $\langle\Delta r^{2}(t)\rangle\sim t$ (line).
 An ensemble generated by cyclic athermal quasi-static shear.
 We apply cyclic athermal quasistatic shear (CAQS) to standard models of soft, frictionless particles in 2D and 3D (see Models and Shear protocol in Methods).
-We present 2D data in the main text and 3D data in the Extended Data.
+We present 2D data in the main text and 3D data in the Supplementary Information (SI).
 For the 2D model, previous studies report a jamming density (J-point density) $\varphi_{\rm J}\approx 0.842-0.843$ O’Hern et al.
@@
 The above results imply that the configurational space is effectively explored by the dynamical trajectory of the system in the irreversible phase, encouraging us to consider a statistical ensemble generated by shear dynamics.
-In the main text below, we fix $\gamma_{\rm max}={0.7}$, and vary $\varphi$ systematically in the window of $[0.833,0.849]$.
+In this study, we fix $\gamma_{\rm max}={0.7}$ unless otherwise specified, and vary $\varphi$ systematically in the window of $[0.833,0.849]$.
 At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ independent configurations are collected to construct the ensemble.
-In the Extended Data, we present additional results obtained for $\gamma_{\rm max}={1}$.
+In the SI, we present additional results obtained for $\gamma_{\rm max}={1}$.
 All configurations are typically isotropic since they are collected at $\gamma=0$.
-Figure 1d shows the time evolution of the coordination number $Z(t)$ in the irreversible phase, obtained from a typical simulation near $\varphi_{\rm J}$.
+Figure 1d shows the evolution of the coordination number $Z(t)$ in the irreversible phase, obtained from a typical simulation near $\varphi_{\rm J}$.
 Rattlers (particles with fewer than $d+1$ contacts) are removed in the computation of $Z$.
 At first glance, one sees the coexistence of jammed ($Z\approx 2d=4$) and unjammed states ($Z=0$), similar to the coexistence of two ferromagnetic states (positive $m$ and negative $m$) in the time evolution of the magnetization $m(t)$ in an Ising model near a first-order phase transition Binder (1987a).
-However, a more careful examination reveals four distinct states, which we discuss below.
-Figure 1: Cyclic athermal quasi-static shear simulations.
-(a) Reversible ($\delta r_{\infty}<0.1$, blue region) and irreversible ($\delta r_{\infty}>0.1$, yellow region) phases ($N={1000}$).
-The solid line represents $\varphi_{\rm J}^{\infty}={0.8432}$, and the dashed line represents $\gamma_{\rm max}=0.7$ used to generate the ensemble in the main text.
-(b) One-cycle displacement $\delta r(t)$ and (d) average coordination number $Z(t)$ for a typical sample.
-(c) The MSD data show diffusive behavior $\langle\Delta r^{2}(t)\rangle\sim t$ (line).
-Data in (b-c) are obtained at a fixed $\varphi=0.841$.
+However, a more careful examination reveals four distinct states, which we discuss in the next section.
+Figure 2: First three shear cycles.
+The (a) strain $\gamma$, (b) sample-averaged coordination number $Z$, (c) sample-averaged pressure $P$, and (d) jamming fraction $F_{\rm J}$ are plotted as functions of $t$ during the first three shear cycles ($\varphi=0.8425$, $N_{\rm s}=768$ samples).
+Orange and blue curves are obtained from rapidly quenched ($\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$) and mechanically trained ($\varphi_{\rm j}\approx 0.848$) initial configurations respectively.
+The jamming fraction $F_{\rm J}$ used in the scaling analysis is obtained at $\gamma=0$ (black cycles in (d)).
+In (e) and (f), plotted is the $Z(t)$ of two typical samples initially generated by rapid quenching.
+The system status at $\gamma=0$ is highlighted by black cycles, which fluctuate among jammed ($Z\geq 4$), fragile ($43<Z<$ 4) and unjammed ($Z=0$) states.
+It is well known now that the jamming density $\varphi_{\rm j}$ is not unique and protocol-dependent Parisi and Zamponi (2010); Chaudhuri et al.
+(2010); Ozawa et al.
+(2012); Kumar and Luding (2016); Kawasaki and Miyazaki (2020); Das et al.
+(2020); Babu et al.
+(2021); Jin and Yoshino (2021) (In this study, $\varphi_{\rm J}$ denotes the J-point density that is the minimum jamming density, and $\varphi_{\rm j}$ denotes the protocol-dependent jamming density): In the compression protocol, $\varphi_{\rm j}$ depends on the compression rate or the density of the initial equilibrium configuration Chaudhuri et al.
+(2010); Ozawa et al.
+(2012); in the cyclic athermal quasistatic compression (CAQC) protocol, $\varphi_{\rm j}$ depends on the volumetric strain amplitude (the maximum density $\varphi_{\rm max}$ to which the system is compressed) and the number $N_{\rm c}$ of cycles Kumar and Luding (2016); Kawasaki and Miyazaki (2020); in the CAQS protocol, $\varphi_{\rm j}$ depends on the shear strain amplitude $\gamma_{\rm max}$ and $N_{\rm c}$ Das et al.
+(2020); Babu et al.
+(2021).
+The continuous range of possible $\varphi_{\rm j}$ is called a jamming-line (J-line), and the minimum jamming density on the J-line is the J-point density $\varphi_{\rm J}$ realized by rapid quenching Babu et al.
+(2021).
+The protocol-dependence of $\varphi_{\rm j}$ in the current model has been investigated in Ref.
+Kawasaki and Miyazaki (2020) using the CAQC protocol, and the maximum jamming density obtained there is $\varphi_{\rm j}\approx 0.8465$ after one over-compression cycle, compared to $\varphi_{\rm J}\approx 0.842-0.843$.
+In order to examine the effects of non-universal $\varphi_{\rm j}$, in this study we prepare two types of samples: rapidly quenched samples with $\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$, and mechanically trained samples generated by the CAQC protocol with $\varphi_{\rm j}\approx 0.848>\varphi_{\rm J}$.
+The protocol-dependence of $\varphi_{\rm j}$ is called a memory effect in Ref.
+Kumar and Luding (2016).
+Below we demonstrate that the cyclic shear simulation employed in this study erases the memory of the initial state: given the initial $\varphi_{\rm j}\approx 0.848>\varphi_{\rm J}$, the jamming density reduces to $\varphi_{\rm J}\approx 0.843$ after one complete shear cycle.
+Figure 2b-d shows the evolution of the sample-averaged coordination number $Z$, the sample-averaged pressure $P$, and the jamming fraction $F_{\rm J}$ of all considered samples at the given $t$.
+The differences between the curves of rapidly quenched ($\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$) and mechanically trained systems ($\varphi_{\rm j}\approx 0.848$) only appear in the first cycle.
+Thus the effect of the variable $\varphi_{\rm j}$ is eliminated during the construction of the ensemble by cyclic shear.
+Previous studies have shown that, independent of the initial jamming density $\varphi_{\rm j}$, if the system is sheared by large strain deformations beyond the yield strain $\gamma_{\rm Y}$, it always evolves into a critical steady state that has a generic jamming density at $\varphi_{\rm J}$, and such a process is accompanied with significant shear dilatancy or hardening effects Babu et al.
+(2021); Pan et al.
+(2023b); Xing et al.
+(2024).
+According to Fig.
+2b-d, the system has indeed reached the steady state before the strain reversal, as the strain amplitude $\gamma_{\rm max}=0.7$ is larger than the yield strain $\gamma_{\rm Y}\sim 0.1$.
+Next we illustrate that the statistical properties of the ensemble are determined by fluctuation effects.
+The overall behavior of the sample-averaged $Z(t)$ and $P(t)$ is very similar to that reported in previous cyclic shear simulations Kumar and Luding (2016); Luding et al.
+(2021): one observes periodic patterns with the sharp onset of unjamming upon strain reversal at integer or half-integer $t$.
+However, no regular patterns can be identified in single-sample curves, which look more like stochastic processes (see Fig.
+2e-f).
+Similar behavior has been observed in previous simple shear simulations Heussinger and Barrat (2009) and experiments Behringer et al.
+(2008), showing that the coexistence of jammed and unjammed states during shear is a generic property of the system, at a fixed $\varphi$ near the jamming transition.
+It is useful to consider two types of fluctuations.
+(i) Fluctuations between different cycles in the simulation of a given sample.
+The configurations at consecutive strain steps $t$ and $t+\delta t$ are related by both affine and non-affine deformations.
+Due to non-affine deformations, the two finite-size configurations can have different structures and jamming densities $\varphi_{\rm J}^{N}$.
+If the preset constant density $\varphi$ in the CAQS simulation is close to the average $\varphi_{\rm J}$, then the two consecutive configurations can be either jammed ($\varphi_{\rm J}^{N}<\varphi$) or unjammed ($\varphi_{\rm J}^{N}>\varphi$).
+This fluctuation effect is the origin of stochastic-like behavior in Fig.
+2e-f.
+(ii) Sample-to-sample fluctuations at a given $t$.
+For similar reasons, at a given $t$, two individual samples can be either jammed or unjammed due to sample-to-sample fluctuations (see marked points at the integer $t$ in Fig.
+2e-f).
+The jamming fraction $F_{\rm J}$ due to such sample-to-sample fluctuations is plotted in Fig.
+2d, whose scaling behavior is analyzed in detail below.
+Note that the fluctuations in (i) and (ii) essentially have the same origin: in finite-size systems, the jamming density $\varphi_{\rm J}^{N}$ has a distribution $p(\varphi_{\rm J}^{N})$ due to the existence of abundant amorphous states.
+As shown in Fig.
+2d, upon on the strain reversal, $F_{\rm J}$ rapidly decays to zero, which means that the configuration jammed along a given shear direction is unstable to reverse shear deformations Kumar and Luding (2016); Boschan et al.
+(2019); Seto et al.
+(2019).
+This reversal-unjamming effect occurs within a strain interval $\delta\gamma\lesssim 0.2$, and after this interval $F_{\rm J}$ reaches a stationary plateau.
+Because the maximum strain $\gamma_{\rm max}>0.2$ in our protocol, the configurations collected in the ensemble are always in the stationary regime.
 Four states: unjammed, jammed, partially crystallized and fragile.
@@
 The red and blue disks are non-rattler and rattler particles respectively.
-We set $\varphi=0.841$ for (a) and (c), $N=256$ for (b) and (c), and $N=1000$ for (e-h).
+We set $\varphi=0.841$ for (a) and (c), $N=256$ for (b) and (c), and $N=1000$ for (a, e-h).
 In Fig.
-2a, we plot the probability distribution $p(Z)$ of the states in the considered ensemble, at a fixed $\varphi=0.841$ in the irreversible phase.
+3a, we plot the probability distribution $p(Z)$ of the states in the considered ensemble, at a fixed $\varphi=0.841$ in the irreversible phase.
 The following four states, represented by peaks in $p(Z)$, can be identified (see Figs.
-2e-h).
+3e-h).
 (i) Unjammed states.
@@
 The delta-peak at $p_{\rm C}(Z)\sim\delta(Z-24/7)$ represents the states with a single unit cell of the hexagonal crystal (see Fig.
-2e).
+3e).
 Here $24/7\approx 3.4286$ is the average number of contacts of the seven particles forming the unit cell.
 Occasionally, states with two or three unit crystal cells can be also found but they are rare.
-(iv) Fragile states. We define the states in the broad peak $43<Z<$ 4, excluding the crystalline peak $p_{\rm C}(Z)$ as the fragile states.
+(iv) Fragile states. We define the states in the broad peak $43<Z<$ 4, excluding the crystalline peak $p_{\rm C}(Z)$, as the fragile states.
 The fractions, $F_{\rm U}$, $F_{\rm J}$, $F_{\rm C}$ and $F_{\rm F}$, of the above four states (i-iv), are obtained by integrating corresponding peaks in $p(Z)$.
 In Fig.
-2b, the fractions are plotted as functions of $\varphi$, showing that $F_{\rm J}(\varphi)$ increases from zero to one across $\varphi_{\rm J}$.
+3b, the fractions are plotted as functions of $\varphi$, showing that $F_{\rm J}(\varphi)$ increases from zero to one across $\varphi_{\rm J}$.
 This behavior is quantitatively similar to $F_{\rm J}(\varphi)$ obtained in previous rapid quench simulations, where finite-size analyses have been carried out to precisely determine the asymptotic jamming transition density $\varphi_{\rm J}^{\infty}$ in the thermodynamic limit $N\rightarrow\infty$ O’hern et al.
@@
 In Fig.
-2c, we report various fractions as functions of $t$ at $\varphi={0.841}$.
+3c, we report various fractions as functions of $t$ at $\varphi={0.841}$.
 The fractions are independent of $t$, confirming that the system is stationary.
@@
 In Fig.
-2d, we plot the maximum coordination numbers, $Z_{\rm J}^{*}$ and $Z_{\rm U}^{*}$, of jammed and fragile peaks.
-Around $\varphi_{\rm J}$, there is a small gap $\delta Z_{\rm gap}\approx 0.2$ between $Z_{\rm J}^{*}$ (upper branch) and $Z_{\rm U}^{*}$ (lower branch).
+3d, we plot the maximum coordination numbers, $Z_{\rm J}^{*}$ and $Z_{\rm F}^{*}$, of jammed and fragile peaks.
+Around $\varphi_{\rm J}$, there is a small gap $\delta Z_{\rm gap}\approx 0.2$ between $Z_{\rm J}^{*}$ (upper branch) and $Z_{\rm F}^{*}$ (lower branch).
 The results are very similar to those obtained by quasistatic uniform shear in Ref.
@@
 Indeed, the fragile states with $43<Z<$ 4 can have percolated force networks of non-rattlers in one or two directions (see Fig.
-2f and g).
+3f and g).
 Thus in our case, anisotropy cannot effectively distinguish fragile from jammed states.
+It should be noted that, due to strong fluctuation effects, the fragile states can be generated either before or after jamming during the shear procedure (see Fig.
+2e-f).
 We demonstrate that the essential difference between fragile and jammed states is their mechanical stability.
 In fragile states, the potential energy per particle is negligibly low (see Stopping criterion for energy minimization in Methods), $\overline{e}_{i}<e_{\rm th}=10^{-20}$ (see Fig.
-3a), but the fraction of non-rattler particles is non-zero.
+4a), but the fraction of non-rattler particles is non-zero.
 These non-rattlers experience forces, which can form a transient, percolated network.
 Such networks are highly heterogeneous (see Fig.
-2f and g), compared to those in jammed states (see Fig.
-2h).
+3f and g), compared to those in jammed states (see Fig.
+3h).
 More importantly, the force networks in fragile states are unstable, revealed by the non-zero net force per particle, $\overline{f}_{i}>f_{\rm th}=10^{-14}$ (see Fig.
-3b).
+4b).
 Thus fragile states are not strictly equilibrated; they turn into unjammed states by sufficiently long relaxation (accurate energy minimization) or mechanical perturbations.
 To demonstrate the instability of fragile states, two tests are carried out.
-First, we perform a compression-decompression perturbation, $\varphi\rightarrow\varphi+\delta\varphi\rightarrow\varphi$, where $\delta\varphi=10^{-8}$, and the energy is minimized after each step.
+First, we perform a compression-decompression perturbation, $\varphi\rightarrow\varphi+\delta\varphi\rightarrow\varphi$, where $\delta\varphi=10^{-8}$, with the energy minimized after each step.
 All fragile states ($43<Z<$ 4) become unjammed ($Z=0$) after this perturbation, while jammed states remain.
 In Fig.
-3c, we plot the distribution $p^{\prime}(Z)$ after the compression-decompression perturbation.
+4c, we plot the distribution $p^{\prime}(Z)$ after the compression-decompression perturbation.
 Independently, without any perturbation we simply regard all fragile states ($43<Z<$ 4) as unjammed ($Z=0$) and recalculate the distribution $\tilde{p}(Z)$.
@@
 The fraction $F_{\rm F}$ of fragile states grows and $F_{\rm U}$ of unjammed states decays with increasing $f_{\rm th}$ (Fig.
-3d), suggesting that many unjammed states become fragile under a looser force-balance condition $\overline{f}_{i}<f_{\rm th}$ with a larger $f_{\rm th}$.
+4d), suggesting that many unjammed states become fragile under a looser force-balance condition $\overline{f}_{i}<f_{\rm th}$ with a larger $f_{\rm th}$.
 In contrast, the constant $F_{\rm J}$ shows that the definition of jammed states is insensitive to the algorithm parameter.
+The $p(Z)$ data with different $f_{\rm th}$ confirm this property: the fragile peak depends on $f_{\rm th}$ while the jammed peak is independent of $f_{\rm th}$ (see SI Fig.
+S1a).
 Note that in granular experiments Bi et al.
@@
 (d) Fractions of four states as functions of $f_{\rm th}$.
-Data are obtained for $N=256$ systems, except for (c).
+Data are obtained for $N=256$ systems at $\varphi=0.841$
+, except for (c).
 Figure 5: Distribution of the coordination number.
@@
 (b) $p(Z)$ at $\varphi=0.843$ for a few different $N$.
-For better visualization, we do not show the unjammed delta-peak at $Z=0$.
+For better visualization, we do not show the unjammed delta-peak at $Z=0$ (see SI Fig.
+S1b for full distributions).
 (c) The fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$.
@@
 Symbols in (b-f) have the same meanings.
-Scaling analysis of $p(Z)$ near the jamming transition.
-The new distribution $p(Z)$ contains two peaks of unjammed and jammed states, whose fractions satisfy $F_{\rm U}(\varphi)+F_{\rm J}(\varphi)=1$ at any $\varphi$.
-In Figs 4a and b, we plot $p(Z)$ for a few different $\varphi$ and $N$, showing that $p(Z)$ is always double-peaked across the jamming transition, which is typical behavior of a first-order rather than a second-order transition.
-To analyze the scaling behavior of $p(Z)$, we consider a general first-order form of first-order phase transitions Binder (1987a, b),
+Scaling analysis near the jamming transition.
+[…]
+(I) Standard first-order transitions without disorder.
+In this case, $F_{1}(x,N)$ has the scaling form $F_{1}(x,N)=\mathcal{F}_{1}[(x-x_{\rm c})N]$.
+The fluctuation-dissipation theorem ensures that the susceptibility can be equivalently computed from the response or from the fluctuations, i.e., $\chi_{\rm dis}=\chi_{\rm con}$.
+Examples in this category include the first-order transition in the $q$-state Potts model with $q>4$ where the control parameter is the temperature $T$, and that in the Ising model driven by an external magnetic field $h$ at temperatures below the critical temperature.
+[…]
+[…]
+In Figs 5a and b, we plot $p(Z)$ for a few different $\varphi$ and $N$, showing that $p(Z)$ is always double-peaked across the jamming transition (see SI Fig.
+S1b), which is typical behavior of a first-order rather than a second-order transition.
+To analyze the scaling behavior of $p(Z)$, we consider a general form of first-order phase transitions Binder (1987a, b),
 $p(Z)=(1-F_{\rm J})\delta(Z)+F_{\rm J}\,p_{\rm J}\left[(Z-Z_{\rm J}^{*})N^{\eta}\right],$ (3)
@@
 The asymptotic jamming density $\varphi_{\rm J}^{\rm\infty}=0.8432$ for $N\rightarrow\infty$ is determined by the intersection of $F_{\rm J}(\varphi)$ curves with different $N$ (see Fig.
-4c), close to the J-point density $\varphi_{\rm J}\approx 0.842-0.843$ reported in previous studies O’hern et al.
+5c).
+This value is consistent with the J-point density $\varphi_{\rm J}\approx 0.842-0.843$ given by previous studies O’hern et al.
 (2003); Vagberg et al.
-(2011).
-The difference between $\varphi_{\rm J}^{\rm\infty}=0.8432$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}=0.8435$ by $\gamma_{\rm max}=1$ (see Extended Data Fig.
-E1) is too small to conclude any systematic dependence on $\gamma_{\rm max}$.
+(2011), which is unambiguously below the maximum protocol-dependent jamming density $\varphi_{\rm j}\approx 0.8465$ reported in Ref.
+Kawasaki and Miyazaki (2020) and $\varphi_{\rm j}\approx 0.848$ obtained by CAQC in this study.
+The difference between $\varphi_{\rm J}^{\rm\infty}=0.8432$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}=0.8435$ by $\gamma_{\rm max}=1$ (see SI Fig.
+S2b and S3) is too small to conclude any systematic dependence on $\gamma_{\rm max}$.
 In Ref.
@@
 As shown in Fig.
-2d, $Z_{\rm J}^{*}$ is independent of $N$ near $\varphi_{\rm J}$, and weakly depends on $\varphi$.
+3d, $Z_{\rm J}^{*}$ is independent of $N$ near $\varphi_{\rm J}$, and weakly depends on $\varphi$.
 In the following scaling analysis, $Z_{\rm J}^{*}$ is approximated by a constant value $Z_{\rm J}^{*}\approx 4.1$ at $\varphi_{\rm J}^{\rm\infty}=0.8432$.
 To keep the expressions concise, we ignore the corrections to the scaling functions from the $\varphi$-dependence of $Z_{\rm J}^{*}$.
-We assume the fraction of jammed states $F_{\rm J}$ having the following scaling form, $F_{\rm J}(\varphi,N)=\mathcal{F}\left[(\varphi-\varphi_{\rm J}^{\infty})N^{\lambda}\right]$.
+We assume the fraction of jammed states $F_{\rm J}$ having the following scaling form, $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}\left[(\varphi-\varphi_{\rm J}^{\infty})N^{\lambda}\right]$.
 The value of the exponent $\lambda$ is important for understanding the nature of the transition.
@@
 Because the free energy is non-singular around a first-order phase transition, it can be expanded, giving $\delta f\sim(T-T_{\rm c})$ to the lowest order.
-Thus, if the jamming transition were a standard first-order phase transition, then $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}\left[(\varphi-\varphi_{\rm J}^{\infty})N\right]$, i.e., $\lambda=1$ (note that $\varphi$ in the granular system is analogous to the temperature $T$ in thermal systems).
+Thus, if the jamming transition were a standard first-order phase transition, then $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}\left[(\varphi-\varphi_{\rm J}^{\infty})N\right]$, i.e., $\lambda=1$ (note that $\varphi$ is the control parameter of the jamming transition).
 However, our numerical data can not be collapsed using $\lambda=1$; in contrast, they can be perfectly collapsed using $\lambda=1/2$ (see Fig.
-4d).
+5d).
 To understand the origin of $\lambda=1/2$ scaling, recall that the athermal jamming transition is not driven by the free energy difference between the two phases.
@@
 (2003): the width of $\rho(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-0.55}$ in both 2D and 3D, supporting our assumption.
-The above assumption also predicts, $\varphi_{\rm J}^{\infty}-\varphi_{\rm J}^{N}\sim N^{-1/2}$, independent of dimensionality.
+The above assumption also predicts, $\varphi_{\rm J}^{\infty}-\varphi_{\rm J}^{N}\sim N^{-1/2}$, independent of the dimensionality.
 The jammed states are defined by $\varphi>\varphi_{\rm J}^{N}$, and thus
 […]
-where $\delta\hat{\varphi}=(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. Equation (4) agrees well with the data in Fig. 4d.
+where $\delta\hat{\varphi}=(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$.
+Equation (4) agrees well with the data in Fig.
+5d.
+The dependence of $F_{\rm J}(\varphi,N)$ on shear parameters is examined in Fig.
+S2.
+The $F_{\rm J}(\varphi,N)$ data are independent of the maximum strain $\gamma_{\rm max}$, and weakly depend on the strain step $\delta\gamma$.
+Independent of $\delta\gamma$, the $F_{\rm J}(\varphi,N)$ data of different $N$ collapse when plotted as a function of the rescaled variable $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$ (see Figs.
+5d and S2c for $\delta\gamma=10^{-2}$ and $10^{-1}$ respectively), showing robustness of the scaling function Eq.
+(4).
 Figure 6: Susceptibilities.
@@
 The solid black lines in (a-c) are Eqs.
-(5), (6) and (7) respectively.
+(5), (6), and the ratio between Eq.
+(5) and the square of Eq.
+(6), respectively.
+The dashed black line in (c) is the second-order expansion Eq.
+(7).
 The dashed red line in (c) is the constant term in Eq.
 (7), i.e., $\chi_{\rm dis}/\chi_{\rm con}^{2}=\pi\sigma_{\varphi}^{2}/2$.
-To draw these lines, we use $u=0.036$ and $\sigma_{\varphi}=0.043$ obtained from the fitting in Fig.
-4d, and $Z_{\rm J}^{*}=4.1$ determined at $\varphi_{\rm J}^{\infty}=0.8432$ in Fig.
-2d.
+To draw these lines, we use $u=0.041$ and $\sigma_{\varphi}=0.043$ obtained from the fitting in Fig.
+5d, and $Z_{\rm J}^{*}=4.1$ determined at $\varphi_{\rm J}^{\infty}=0.8432$ in Fig.
+3d.
 No fitting is performed here.
@@
 (8).
-The open squares in (d) are $\chi_{\rm J}$ data obtained in a small pressure window $0.00110.0009<P<$ 0.0011, corresponding to $\varphi\approx 0.843$.
+The open squares in (d) are $\chi_{\rm J}$ data obtained in a small pressure window $0.00110.0009<P<$ 0.0011, corresponding to $\Delta Z\approx 0.15$ ($\varphi\approx 0.843-0.844$).
 The fitting parameter $\mu$ is plotted in (f) as a function of $\varphi$, where the error bar represents the fitting error.
-[…]
+Once the scaling behavior of $F_{\rm J}(\varphi,N)$ is obtained, it is easy to derive scaling forms of susceptibilities. From Eq. (3) and the definition of $\chi_{\rm dis}$, we obtain,
 $\frac{\chi_{\rm dis}(\varphi,N)}{N}\approx(Z_{\rm J}^{*})^{2}[1-F_{\rm J}(\varphi,N)]F_{\rm J}(\varphi,N),$ (5)
@@
 Equation (5) suggests a scaling form $\chi_{\rm dis}(\varphi,N)/N=\mathcal{X}_{\rm dis}\left[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}\right]$, in agreement with the data in Fig.
-5a.
-The alternative connected susceptibility is defined as $\chi_{\rm con}=d\langle Z\rangle/d\varphi$.
+6a.
 From Eq.
 (3), one finds that $\langle Z\rangle\approx F_{\rm J}Z_{\rm J}^{*}$ and thus $\langle Z\rangle$ has the same scaling form as $F_{\rm J}$, i.e., $\langle Z\rangle(\varphi,N)=\mathcal{Z}[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}]$, consistent with the data in Fig.
-4e and f.
+5e and 5f.
 Using $\chi_{\rm con}=Z_{\rm J}^{*}\frac{dF_{\rm J}}{d\varphi}$ and Eq.
-(4), we derive,
+(4), we obtain,
 $\frac{\chi_{\rm con}(\varphi,N)}{N^{1/2}}\approx\frac{Z_{\rm J}^{*}}{\sigma_{\varphi}\sqrt{2\pi}}\exp\left[-\left(\frac{\delta\hat{\varphi}+u}{\sqrt{2}\sigma_{\varphi}}\right)^{2}\right],$ (6)
@@
 (7) gives a scaling, $\chi_{\rm dis}\sim\chi_{\rm con}^{2},$ which is a key signature of the presence of quenched disorder.
-This square relationship appears generally at first-order transitions in disordered systems, such as the first-order phase transition in the RFIM Nattermann (1998), the brittle yielding of amorphous solids Ozawa et al.
-(2018) and the melting of ultrastable hard-sphere glasses Zhang et al.
-(2022).
 Comparing to the simulation data, Eq.
 (7) works well around the extreme point (see Fig.
-5c).
+6c).
+The agreement can be improved using higher-order expansions.
 It seems that the non-Gaussian effect in the distribution $\rho(\varphi_{\rm J}^{N})$, which is neglected in the present analysis, is amplified in the data of $\chi_{\rm dis}/\chi_{\rm con}^{2}$, resulting in slight asymmetry.
-Extended Data Figs.
-E1-E4 show that the scaling functions Eqs.
+In SI Figs.
+S3-S6, we show that the scaling functions Eqs.
 (4)-(7) work in 3D, and for a different max strain $\gamma_{\max}=1$ in 2D.
-Finally let us discuss the finite-size effects of the jammed peak $p_{\rm J}(Z)$ in Eq.
+In 3D, we obtain $\varphi_{\rm J}^{\infty}=0.648$, consistent with the previously reported J-point density $\varphi_{\rm J}\approx 0.648$ of this model Das et al.
+(2020), which is the minimum jamming density on the J-line.
+For the same 3D model, the protocol-dependent jamming density is obtained up to $\varphi_{\rm j}=0.6616$ by the thermal annealing method Chaudhuri et al.
+(2010), and up to $\varphi_{\rm j}=0.661$ by the CAQS mechanical training method Das et al.
+(2020).
+Next let us discuss the finite-size effects of the jammed peak $p_{\rm J}(Z)$ in Eq.
 (3).
 For standard first-order phase transitions, $\eta=1/2$ Binder (1987b, a).
-Thus in that scenario the susceptibility of the jamming peak would scale as, $\chi_{\rm J}\equiv N\sigma_{\rm J}^{2}=N\langle(Z-Z_{\rm J}^{*})^{2}\rangle_{\rm J}\sim N^{0}$, where $\sigma_{\rm J}^{2}$ is the variance of $p_{\rm J}(Z)$ and $\langle\ldots\rangle_{\rm J}$ represents the average restricted to the jammed states only.
+Thus in that scenario the susceptibility of the jamming peak would scale as, $\chi_{\rm J}\equiv N\sigma_{\rm J}^{2}=N\langle(Z-Z_{\rm J}^{*})^{2}\rangle_{\rm J}\sim{\rm constant}$, where $\sigma_{\rm J}^{2}$ is the variance of $p_{\rm J}(Z)$ and $\langle\ldots\rangle_{\rm J}$ represents the average restricted to the jammed states only.
 However, our $\chi_{\rm J}$ data disagree with this scaling (see Figs.
-5d-f).
-At different $\varphi$, $\chi_{\rm J}$ can be fitted to a pow-law form,
+6d-f).
+At different $\varphi$, $\chi_{\rm J}$ can be fitted to a pow-law form $\chi_{\rm J}\sim\ell^{-\mu}$ Hexner et al.
+(2018, 2019), or
 $\chi_{\rm J}\sim N^{-\mu/d},$ (8)
-where $\mu=d(2\eta-1)$ is a fitting parameter.
+where $N\sim\ell^{d}$.
+The exponent $\mu=d(2\eta-1)$ is a fitting parameter.
 At large or small $\varphi$ away from $\varphi_{\rm J}^{\rm\infty}$ (see Fig.
-5f), the exponent is close to zero ($\mu\approx-0.2$), suggesting that the local coordination numbers $Z_{i}$ are uniformly distributed.
+6f), the exponent is close to zero ($\mu\approx-0.2$), suggesting that the local coordination numbers $Z_{i}$ are uniformly distributed.
 The deviation from the uncorrelated behavior ($\mu=0$) is significant around $\varphi_{\rm J}^{\rm\infty}$, where $\mu\approx-0.8$ reaches the minimum.
@@
 However, the negative $\mu$ in Fig.
-5f is not due to the criticality of the jamming transition.
+6f is not due to the criticality of the jamming transition.
 Here it is essential to consider the volume fluctuations.
@@
 If we instead select configurations around a constant pressure $P$, then the fluctuations become significantly smaller, and $\mu\approx 0$ (Fig.
-5d).
+6d).
 This observation is consistent with a previous study Ikeda (2023), suggesting that the large fluctuation of $Z$ near $\varphi_{\rm J}^{\rm\infty}$ in the $\varphi$-controlled protocol might be originated from the fluctuation of $\varphi_{\rm J}^{N}$.
@@
 (2018, 2019).
+Ref.
+Hexner et al.
+(2019) also compares an ensemble of subsystems cut out from large packings, to an ensemble of whole systems with the same volume under periodic boundary conditions (similar to the $P$-controlled ensemble considered in this study): the contact fluctuations in the former are larger than those in the latter, and the convergence of these two are expected to occur in extremely large systems that are not attainable in current simulations – Due to this pre-asymptotic effect, the hyperuniform exponent $\mu\simeq 1$ is only observable in the first ensemble, while in the second ensemble, the apparent exponent is close to $\mu\approx 0$, consistent with our $P$-controlled data in Fig.
+6d.
+In short, although the hyperuniform exponent $\mu=1$ is not directly observed, our results do not contradict the recently reported hyperuniform behavior of contact distributions at jamming.
 We emphasize that the finite-size effects of the jammed peak $p_{\rm J}(Z)$ contribute negligibly to the scaling of $F(\varphi,N)$, $\chi_{\rm dis}(\varphi,N)$ and $\chi_{\rm con}(\varphi,N)$.
 In other words, the first-order nature of the jamming transition is independent of how $Z_{i}$ is spatially distributed in jammed packings.
+In the scaling analysis presented above, $Z$ is treated as the essential order parameter.
+In principle, the same kind of scaling analysis can be applied to any physical quantity $A$ that varies discontinuously at jamming (e.g., $A$ can be the bulk modulus $B$ or the fraction of non-rattlers $f_{\rm NR}$).
+The distribution $p(A)$ should have the same form as Eq.
+(3), $p(A)=(1-F_{\rm J})\delta(A)+F_{\rm J}p_{\rm J}(A)$.
+With fragile states removed, generally $A=0$ below jamming because no contacts are formed (e.g., $B=0$ and $f_{\rm NR}=0$), and thus the unjammed peak is always a delta function.
+The fraction of jammed states $F_{\rm J}$ only depends on the constructed ensemble, which determines how the states are sampled, and thus its scaling form $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}]$ is independent of $A$.
+The jammed peak $p_{\rm J}(A)$ does depend on $A$, but as shown above, $p_{\rm J}(A)$ is irrelevant to the interested scaling behavior of $F_{\rm J},\chi_{\rm dis}$ and $\chi_{\rm con}$.
+Thus, our conclusion drawn from the scaling analysis should be robust and generic, independent of which parameter $A$ is chosen for the scaling analysis, as long as $A$ can reflect the discontinuous nature of jamming.
 Conclusion.
-In this study, we understand the nature of jamming transition through an ensemble approach analogous to the statistical mechanics in equilibrium systems Baule et al.
+In this study, we investigate the nature of the jamming transition through an ensemble approach analogous to the statistical mechanics in equilibrium systems Baule et al.
 (2018); Bi et al.
 (2015).
-Within such a framework, the order of the jamming transition is independent of the complex properties of jammed packings, including isostaticity, marginality and hyperuniformity (see (i-iii) in Introduction).
-As a first-order jamming transition, the finite-size scaling exponents measured in this study are naturally universal in 2D and 3D, without the need to introduce the upper critical dimension $d_{\rm u}=2$ Parisi et al.
-(2018), a concept relevant to continuous transitions only.
+Within such a framework, jamming is demonstrated to be a first-order transition with quenched disorder, in the quasistatic deformation limit where all fragile states are excluded.
+The order of the jamming transition is independent of the complex properties of jammed packings, including isostaticity, marginality and hyperuniformity (see (i-iii) in Introduction).
+In previous rapid quench simulations, it is observed that the width $w$ of the jamming density distribution $P(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-\Omega}$ with $\Omega=0.55\pm 0.03$ in both 2D and 3D O’hern et al.
+(2003).
+This finite-size scaling depends on the total number of particles, $N$, rather than on the system length $\ell$, and the exponent $\Omega\approx 1/2$ is independent of the dimensionality within the numerical accuracy.
+An explanation, as suggested previously, is that jamming is a second-order transition with an upper critical dimension $d_{\rm u}=2$ Goodrich et al.
+(2012).
+Here we propose an alternative interpretation: jamming is a first-order transition with quenched disorder, and consequently the finite-size scaling is a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$, independent of the dimensionality.
 It is of particular interest to reconcile the first-order transition established here under quasi-static shearing and the second-order transition observed in previous finite-rate rheology (see (iv) in Introduction).
-The jamming transition appears to be a rare example with the simultaneous existence of first-order and second-order transitions between liquid and solid states.
-Conventionally, the two kinds of transitions only coexist in gas-liquid systems, but not in liquid-crystalline solid systems.
+Conventionally, first- and second-order transitions coexist in gas-liquid systems, but not in liquid-crystalline solid systems.
+Because the liquid-crystalline solid transition is accompanied with spontaneous symmetry breaking, it cannot be a continuous type.
+However, there is no apparent spontaneous symmetry breaking during liquid-amorphous solid transitions (such as the jamming transition).
+Thus there is no fundamental reason to exclude the possibility of the coexistence of discontinuous and continuous transitions between a liquid and an amorphous solid.
+This study opens an avenue to unify first-order (this study) and second-order Olsson and Teitel (2007) jamming transitions.
+Another direction of the future work is to study the competition between the sample-to-sample fluctuations due to disorder and the fluctuations associated to the granular temperature Pöschel and Luding (2001); Luding et al.
+(2021), and examine the “thermal effects” on the nature of the first-order jamming transition.
+Similar studies on the thermal vestiges of the zero-temperature physics have been recently carried out in the driven RFIM at finite temperatures Yao and Jack (2023).
+It would be also interesting to examine if friction can alter the order of the jamming transition Song et al.
+(2008); Ramaswamy et al.
+(2023).
 Methods.
@@
 The system is then rapidly quench compressed to the target $\varphi$.
-The CAQS is performed under the Lees-Edwards boundary conditions, with a fixed $\varphi$.
+The CAQS is performed under the Lees-Edwards boundary conditions Lees and Edwards (1972), with a fixed $\varphi$.
+The strain $\gamma$ is varied stepwise, rather than continuously, between 0 and $\gamma_{\rm max}$.
 We use a strain step $\delta\gamma=0.1$ to generate the phase diagram in Fig.
@@
 During a cycle, the strain is varied as $\gamma=0\rightarrow\gamma_{\rm max}\rightarrow 0$.
-We set $\gamma_{\rm max}=0.7$ in the main text and $\gamma_{\rm max}=1$ in the Extended Data for the 2D model, and $\gamma_{\rm max}=0.5$ for the 3D model (Extended Data).
+We set $\gamma_{\rm max}=0.7$ in the main text and $\gamma_{\rm max}=1$ in the SI for the 2D model, and $\gamma_{\rm max}=0.5$ for the 3D model (SI).
 The number of cycles is represented by $t$ with unit oscillation period ($t=1$).
@@
 Acknowledgements.
-We warmly thank Hajime Yoshino for inspiring discussions.
+We warmly thank Hajime Yoshino, Stephen Teitel, Bulbul Chakraborty, Olivier Dauchot, Gilles Tarjus for inspiring discussions.
 We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 11974361 and 12047503), from Chinese Academy of Sciences (Grants ZDBS-LY-7017 and KGFZD-145-22-13), and from Wenzhou Institute (Grant WIUCASQD2023009).
 In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
+Supplementary Information
+Figure S1: Additional data on the distribution $p(Z)$ in two dimensions.
+(a) $p(Z)$ for a few different $f_{\rm th}$ ($\varphi=0.841$ and $N=1000$).
+(b) Full probability distribution $\tilde{p}(Z)$ at a few different $\varphi$ near $\varphi_{\rm J}^{\infty}=0.8432$ ($N=1000$ and $f_{\rm th}=10^{-14}$).
+Figure S2: Additional data on the jamming fraction $F_{\rm J}$ in two dimensions.
+(a) $F_{\rm J}$ as a function of $\varphi$, for different strain steps $\delta\gamma$, with fixed $\gamma_{\rm max}=0.7$ and $N=512$.
+(b) $F_{\rm J}$ as a function of $\varphi$, for different maximum strains $\gamma_{\rm max}$, with fixed $\delta\gamma=10^{-2}$ and $N=512$.
+(c) Data collapse of $F_{\rm J}(\varphi,N)$ as a function of rescaled variable $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$, with $\varphi_{\rm J}^{\infty}=0.8432$, $\delta\gamma=10^{-1}$ and $\gamma_{\rm max}=0.7$.
+Figure S3: Distribution of the coordination number in two dimensions for $\gamma_{\rm max}=1$.
+(a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$.
+The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8435$.
+(b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$.
+The solid line represents the fitting according to Eq.
+(4),with two fitting parameters $u=0.051$ and $\sigma_{\varphi}=0.043$.
+Figure S4: Susceptibilities in two dimensions for $\gamma_{\rm max}=1$.
+(a) The disconnected susceptibility $\chi_{\rm dis}$ rescaled by $N$, (b) the connected susceptibility $\chi_{\rm con}$ rescaled by $N^{1/2}$, and (c) the ratio between $\chi_{\rm dis}$ and $\chi_{\rm con}^{2}$ are plotted as functions of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$.
+The solid black lines in (a-c) are Eqs.
+(5), (6) and (7) respectively.
+The dashed red line in (c) is the constant term in Eq.
+(7), i.e., $\chi_{\rm dis}/\chi_{\rm con}^{2}=\pi\sigma_{\varphi}^{2}/2$.
+To draw these lines, we use $u=0.051$ and $\sigma_{\varphi}=0.043$ obtained from the fitting in Fig.
+S3b.
+Symbols in (a-c) have the same meanings as indicated in the legend of (a).
+Figure S5: Distribution of the coordination number in three dimensions.
+(a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$.
+The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.648$.
+(b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm j}^{\infty})N^{1/2}$.
+The solid line represents the fitting according to Eq.
+(4), with two fitting parameters $u=0.077$ and $\sigma_{\varphi}=0.051$.
+Figure S6: Susceptibilities in three dimensions.
+(a) The disconnected susceptibility $\chi_{\rm dis}$ rescaled by $N$, (b) the connected susceptibility $\chi_{\rm con}$ rescaled by $N^{1/2}$, and (c) the ratio between $\chi_{\rm dis}$ and $\chi_{\rm con}^{2}$ are plotted as functions of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$.
+The solid black lines in (a-c) are Eqs.
+(5), (6) and (7) respectively.
+The dashed red line in (c) is the constant term in Eq.
+(7), i.e., $\chi_{\rm dis}/\chi_{\rm con}^{2}=\pi\sigma_{\varphi}^{2}/2$.
+To draw these lines, we use $u=0.077$ and $\sigma_{\varphi}=0.051$ obtained from the fitting in Fig.
+S5b.
+Symbols in (a-c) have the same meanings as indicated in the legend of (a).

### v4

@@
 Based on scaling analyses, we show that fluctuations of the jamming density in finite-sized systems have important consequences on the finite-size effects of various quantities, resulting in a square relationship between disconnected and connected susceptibilities, a key signature of the first-order transition with quenched disorder.
-This study puts the jamming transition into the category of a broad class of transitions in disordered systems where sample-to-sample fluctuations dominate over thermal fluctuations, suggesting that the nature and behavior of the jamming transition can be understood within the developed theoretical framework of the athermally driven random-field Ising model.
-Introduction.
+This study puts the jamming transition into the category of a broad class of transitions in disordered systems where sample-to-sample fluctuations dominate over thermal fluctuations, suggesting that the nature and behavior of the jamming transition might be better understood within the developed theoretical framework of the athermally driven random-field Ising model.
+Introduction
 Jamming in athermal particles is a paradigm of transitions between fluids and amorphous solids Makse et al.
@@
 (2005), where $\Delta Z=Z-Z_{\rm iso}$ is the excess coordination number.
-The effective medium theory gives a scattering length scale diverging at $\varphi_{\rm J}$, $l_{c}\sim\Delta Z^{-1/2}\sim(\varphi-\varphi_{\rm J})^{-1/4}$, below which continuum elasticity breaks down Wyart (2010); DeGiuli et al.
+The effective medium theory gives a scattering length scale diverging at $\varphi_{\rm J}$, $l_{\rm c}\sim\Delta Z^{-1/2}\sim(\varphi-\varphi_{\rm J})^{-1/4}$, below which continuum elasticity breaks down Wyart (2010); DeGiuli et al.
 (2014a); Lerner et al.
 (2014).
-Other related diverging length scales include the transverse wavelength, $\xi_{T}\sim(\varphi-\varphi_{\rm J})^{-\nu_{T}}$ with $\nu_{T}\approx 0.24$ Silbert et al.
-(2005), and the longitudinal wavelength, $\xi_{L}\sim(\varphi-\varphi_{\rm J})^{-\nu_{L}}$ with $\nu_{L}\approx 0.48$ Silbert et al.
+Other related diverging length scales include the transverse wavelength, $\xi_{\rm T}\sim(\varphi-\varphi_{\rm J})^{-\nu_{\rm T}}$ with $\nu_{\rm T}\approx 0.24$ Silbert et al.
+(2005), and the longitudinal wavelength, $\xi_{\rm L}\sim(\varphi-\varphi_{\rm J})^{-\nu_{\rm L}}$ with $\nu_{\rm L}\approx 0.48$ Silbert et al.
 (2005).
@@
 In other words, the caging correlation length, $\xi_{\rm G}\sim\infty$, remains infinite near jamming, when the transition is approached ($\varphi\rightarrow\varphi_{\rm J}$) from below ($\varphi<\varphi_{\rm J}$).
-The mean-field replica theory predicts the values of exponents in the weak force and small gap distributions, $\alpha=0.41269$ and $\theta=0.42311$, coinciding with the relationship $\theta=1/\alpha-2$ given by the mechanical marginal stability analysis Charbonneau et al.
+The mean-field replica theory predicts the values of exponents in the weak force and small gap distributions, $\alpha=0.41\ldots$ and $\theta=0.42\ldots$, coinciding with the relationship $\theta=1/\alpha-2$ given by the mechanical marginal stability analysis Charbonneau et al.
 (2014, 2015b); Wang et al.
 (2022).
-The theory provides an additional scaling relationship between the cage size $\Delta$ and the entropic pressure $p$, $\Delta\sim p^{-\kappa}$ with $\kappa=1.41574$.
-The exponents appear to be independent of the dimensionality $d$ for $d\geq d_{\rm u}$ Charbonneau et al.
+The theory provides an additional scaling relationship between the cage size $\Delta$ and the entropic pressure $p$, $\Delta\sim p^{-\kappa}$ with $\kappa=1.41\ldots$ The exponents appear to be independent of the dimensionality $d$ for $d\geq d_{\rm u}$ Charbonneau et al.
 (2014, 2015b); Parisi et al.
@@
 Apparently, as the signature of a first-order transition, this discontinuity is inconsistent with the viewpoint of a continuous transition described above.
-Note that not all physical quantities exhibit a discontinuity at jamming: for example, the pressure $P$ and the shear modulus $G$ vanish continuously when the jamming transition is approached from above Makse et al.
+Note that not all physical quantities exhibit a discontinuity at jamming: some quantities, such as the pressure $P$ and the shear modulus $G$, vanish continuously when the jamming transition is approached from above Makse et al.
 (2000); O’Hern et al.
-(2003).
+(2003) (for example, see Ref.
+Kumar and Luding (2016)) for the relation between $P$ and $\varphi$).
 On the other hand, other quantities, such as the bulk modulus $B$ and the fraction of non-rattlers $f_{\rm NR}$, jump abruptly at jamming, similar to $Z$.
@@
 (2003).
+The protocol dependence of the jamming density is not observed in our simulations: the preparation history of the initial condition becomes irrelevant after the first shear cycle, because a steady state is reached before the strain reversal.
 Third, in our ensemble, we carefully exclude partially crystallized and fragile states with $Z<Z_{\rm iso}$ that are sensitive to mechanical perturbations or protocol parameters, and regard them as unjammed states with $Z=0$.
+Any transient and intermediate states generated during shear, which appear before the steady states, are excluded in our analysis.
 If such states are included, the discontinuity in $Z$ diminishes and the jamming transition looks continuous, similar to the results reported in Ref.
@@
 (inset) The MSD data show diffusive behavior $\langle\Delta r^{2}(t)\rangle\sim t$ (line).
+Results
 An ensemble generated by cyclic athermal quasi-static shear.
@@
 In this study, we fix $\gamma_{\rm max}={0.7}$ unless otherwise specified, and vary $\varphi$ systematically in the window of $[0.833,0.849]$.
-At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ independent configurations are collected to construct the ensemble.
+At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ independent configurations are generated to construct the ensemble.
 In the SI, we present additional results obtained for $\gamma_{\rm max}={1}$.
-All configurations are typically isotropic since they are collected at $\gamma=0$.
+All configurations are collected at $\gamma=0$, which have weak anisotropy characterized by a non-zero average macroscopic friction coefficient (the ratio between stress and pressure), $\langle\sigma/P\rangle\approx 0.08$.
 Figure 1d shows the evolution of the coordination number $Z(t)$ in the irreversible phase, obtained from a typical simulation near $\varphi_{\rm J}$.
@@
 Kawasaki and Miyazaki (2020) using the CAQC protocol, and the maximum jamming density obtained there is $\varphi_{\rm j}\approx 0.8465$ after one over-compression cycle, compared to $\varphi_{\rm J}\approx 0.842-0.843$.
-In order to examine the effects of non-universal $\varphi_{\rm j}$, in this study we prepare two types of samples: rapidly quenched samples with $\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$, and mechanically trained samples generated by the CAQC protocol with $\varphi_{\rm j}\approx 0.848>\varphi_{\rm J}$.
+In order to examine the effects of non-universal $\varphi_{\rm j}$, in this study we prepare two types of samples: rapidly quenched samples with $\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$, and mechanically trained samples generated by the CAQC protocol with $\varphi_{\rm j}={0.8479(4)}>\varphi_{\rm J}$.
 The protocol-dependence of $\varphi_{\rm j}$ is called a memory effect in Ref.
@@
 It was proposed that fragile and jammed states differ in the percolation of the strong force network: in fragile states the percolation occurs anisotropically in one direction only, while in jammed states it occurs isotropically in both directions.
-In this study, all states are collected during cyclic shear at $\gamma=0$, without systematically introducing anisotropy.
-Indeed, the fragile states with $43<Z<$ 4 can have percolated force networks of non-rattlers in one or two directions (see Fig.
+In this study, all states are collected during cyclic shear at $\gamma=0$, and the fragile states with $43<Z<$ 4 can have percolated force networks of non-rattlers in one or two directions (see Fig.
 3f and g).
@@
 It provides a protocol to examine our scenario of fragile states in future experimental studies.
-Because fragile states are unstable, from now on we count them as unjammed states.
+Because fragile (and partially crystallized) states with $43<Z<$ 4 are unstable, from now on we count them as unjammed states.
 More specifically, we replace the original distribution $p(Z)$ with the modified distribution $\tilde{p}(Z)$.
@@
 (d) Fractions of four states as functions of $f_{\rm th}$.
-Data are obtained for $N=256$ systems at $\varphi=0.841$
-, except for (c).
+Data are obtained for $N=256$ systems at $\varphi=0.841$, except for (c).
 Figure 5: Distribution of the coordination number.
@@
 (c) The fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$.
-The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8432$.
+The intersection of curves gives $\varphi_{\rm J}^{\infty}={0.8432(2)}$.
 (d) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$.
 The solid line represents fitting to Eq.
-(4), with two fitting parameters $u=0.041$ and $\sigma_{\varphi}=0.043$.
+(4), with two fitting parameters $u={0.041(1)}$ and $\sigma_{\varphi}={0.043(1)}$.
 The average coordination number $\langle Z\rangle$ is plotted as a function of $\varphi$ in (e) and of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$ in (f).
@@
 where $\delta(Z)$ and $p_{\rm J}(Z)$ correspond to the unjammed and jammed peaks respectively.
-The asymptotic jamming density $\varphi_{\rm J}^{\rm\infty}=0.8432$ for $N\rightarrow\infty$ is determined by the intersection of $F_{\rm J}(\varphi)$ curves with different $N$ (see Fig.
+The asymptotic jamming density $\varphi_{\rm J}^{\rm\infty}={0.8432(2)}$ for $N\rightarrow\infty$ is determined by the intersection of $F_{\rm J}(\varphi)$ curves with different $N$ (see Fig.
 5c).
@@
 Kawasaki and Miyazaki (2020) and $\varphi_{\rm j}\approx 0.848$ obtained by CAQC in this study.
-The difference between $\varphi_{\rm J}^{\rm\infty}=0.8432$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}=0.8435$ by $\gamma_{\rm max}=1$ (see SI Fig.
-S2b and S3) is too small to conclude any systematic dependence on $\gamma_{\rm max}$.
+The difference between $\varphi_{\rm J}^{\rm\infty}={0.8432(2)}$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}={0.8435(1)}$ by $\gamma_{\rm max}=1$ (see SI Fig.
+S2b and S3-4) is too small to conclude any systematic dependence on $\gamma_{\rm max}$.
 In Ref.
@@
 (2021).
+We emphasize that $\gamma_{\rm max}=0.7$ in this study is significantly larger than the yield strain $\gamma_{\rm Y}\approx 0.1$ (see SI Fig.
+S5a).
+For a small strain amplitude, $\gamma_{\rm max}=0.05$, we observe a logarithmic growth of $\varphi$ as a function of $t$ under a constant pressure condition Knight et al.
+(1995); Kumar and Luding (2016), suggesting that the jamming density can be increased by cyclic shear with a small $\gamma_{\rm max}$ Das et al.
+(2020); Babu et al.
+(2021); in contrast, for $\gamma_{\rm max}=0.7$, $\varphi$ remains as a constant (see SI Fig.
+S5b)
 As shown in Fig.
 3d, $Z_{\rm J}^{*}$ is independent of $N$ near $\varphi_{\rm J}$, and weakly depends on $\varphi$.
-In the following scaling analysis, $Z_{\rm J}^{*}$ is approximated by a constant value $Z_{\rm J}^{*}\approx 4.1$ at $\varphi_{\rm J}^{\rm\infty}=0.8432$.
+In the following scaling analysis, $Z_{\rm J}^{*}$ is approximated by a constant value $Z_{\rm J}^{*}{=4.1(1)}$ at $\varphi_{\rm J}^{\rm\infty}=0.8432$.
 To keep the expressions concise, we ignore the corrections to the scaling functions from the $\varphi$-dependence of $Z_{\rm J}^{*}$.
@@
 Note that $\rho(\varphi_{\rm J}^{N})$ has been explicitly measured in the compression protocol O’Hern et al.
-(2002); O’hern et al.
-(2003): the width of $\rho(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-0.55}$ in both 2D and 3D, supporting our assumption.
+(2002, 2003): the width of $\rho(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-0.55}$ in both 2D and 3D, supporting our assumption.
 The above assumption also predicts, $\varphi_{\rm J}^{\infty}-\varphi_{\rm J}^{N}\sim N^{-1/2}$, independent of the dimensionality.
@@
 In SI Figs.
-S3-S6, we show that the scaling functions Eqs.
+S3-4 and S6-7, we show that the scaling functions Eqs.
 (4)-(7) work in 3D, and for a different max strain $\gamma_{\max}=1$ in 2D.
-In 3D, we obtain $\varphi_{\rm J}^{\infty}=0.648$, consistent with the previously reported J-point density $\varphi_{\rm J}\approx 0.648$ of this model Das et al.
+In 3D, we obtain $\varphi_{\rm J}^{\infty}={0.6479(2)}$, consistent with the previously reported J-point density $\varphi_{\rm J}\approx 0.648$ of this model Das et al.
 (2020), which is the minimum jamming density on the J-line.
@@
 Thus, our conclusion drawn from the scaling analysis should be robust and generic, independent of which parameter $A$ is chosen for the scaling analysis, as long as $A$ can reflect the discontinuous nature of jamming.
-Conclusion.
+If one instead chooses a continuous variable, such as the pressure $P$, then the above scaling analysis would not work.
+Discussion
 In this study, we investigate the nature of the jamming transition through an ensemble approach analogous to the statistical mechanics in equilibrium systems Baule et al.
@@
 Here we propose an alternative interpretation: jamming is a first-order transition with quenched disorder, and consequently the finite-size scaling is a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$, independent of the dimensionality.
+We expect that the same finite-size scaling applies to the steady state configurations generated by uniform simple shear.
 It is of particular interest to reconcile the first-order transition established here under quasi-static shearing and the second-order transition observed in previous finite-rate rheology (see (iv) in Introduction).
@@
 This study opens an avenue to unify first-order (this study) and second-order Olsson and Teitel (2007) jamming transitions.
-Another direction of the future work is to study the competition between the sample-to-sample fluctuations due to disorder and the fluctuations associated to the granular temperature Pöschel and Luding (2001); Luding et al.
+In this study, we ignore thermal fluctuations.
+An interesting direction of the future work is to study the competition between the sample-to-sample fluctuations due to disorder and the fluctuations associated to the granular temperature Pöschel and Luding (2001); Luding et al.
 (2021), and examine the “thermal effects” on the nature of the first-order jamming transition.
@@
 (2023).
-Methods.
+Methods
 Models.
@@
 where $\epsilon=1$ is the energy unit, $r_{ij}$ the distance between particles $i$ and $j$, $\sigma_{ij}=\frac{\sigma_{i}+\sigma_{j}}{2}$ the mean diameter, and $\Theta(x)$ the Heaviside step function.
-We set unit particle mass, and the diameter of small particles as unit length.
+We set unit particle mass, and the mean diameter as the unit length.
 Shear protocol.
@@
 We set $e_{\rm th}=10^{-20}$ and $f_{\rm th}=10^{-14}$ unless otherwise specified.
-Acknowledgements.
-We warmly thank Hajime Yoshino, Stephen Teitel, Bulbul Chakraborty, Olivier Dauchot, Gilles Tarjus for inspiring discussions.
+Data availability
+Source data are provided with this paper.
+Code availability
+The simulation program is available at https://github.com/dengpan-cn/Library-of-Athermal-Quasi-static-Shear-Simulation-of-Granular-Matter.
+Acknowledgments
+We warmly thank Hajime Yoshino, Stephen Teitel, Bulbul Chakraborty, Olivier Dauchot, Gilles Tarjus, Yujie Wang for inspiring discussions.
 We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 11974361 and 12047503), from Chinese Academy of Sciences (Grants ZDBS-LY-7017 and KGFZD-145-22-13), and from Wenzhou Institute (Grant WIUCASQD2023009).
 In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
+Author Contributions Statement
+All authors contributed equally to this work.
+Competing Interests Statement
+The authors declare no competing interests.
 Supplementary Information
@@
 (c) Data collapse of $F_{\rm J}(\varphi,N)$ as a function of rescaled variable $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$, with $\varphi_{\rm J}^{\infty}=0.8432$, $\delta\gamma=10^{-1}$ and $\gamma_{\rm max}=0.7$.
-Figure S3: Distribution of the coordination number in two dimensions for $\gamma_{\rm max}=1$.
+Figure S3: Jamming fraction in two dimensions for $\gamma_{\rm max}=1$.
 (a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$.
-The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8435$.
+The intersection of curves gives $\varphi_{\rm J}^{\infty}={0.8435(1)}$.
 (b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$.
 The solid line represents the fitting according to Eq.
-(4),with two fitting parameters $u=0.051$ and $\sigma_{\varphi}=0.043$.
+(4),with two fitting parameters $u={0.051(1)}$ and $\sigma_{\varphi}={0.043(1)}$.
 Figure S4: Susceptibilities in two dimensions for $\gamma_{\rm max}=1$.
@@
 Symbols in (a-c) have the same meanings as indicated in the legend of (a).
-Figure S5: Distribution of the coordination number in three dimensions.
+Figure S5: Constant-pressure shear simulations in two dimensions.
+(a) Stress $\sigma$ as a function of the strain $\gamma$ in simple athermal quasistatic shear, under a constant pressure $P=10^{-5}$.
+The yield strain is $\gamma_{\rm Y}\approx 0.1$ (arrow), beyond which the curve becomes nearly a plateau.
+(b) $\varphi$ as a function of the number of cycles $t$ in cyclic athermal quasistatic shear (in one cycle $\gamma$ varies as $0\rightarrow\gamma_{\rm max}\rightarrow-\gamma_{\rm max}\rightarrow 0$), with a fixed $P=10^{-5}$.
+The data of $\gamma_{\rm max}=0.05$ are fitted to a logarithmic form (dashed line), $\varphi(t)=\varphi_{\rm f}-\frac{\varphi_{\rm f}-\varphi_{0}}{1+B\log(1+t/\tau)}$, as proposed in Ref.
+[74].
+Here $\varphi_{0}=0.8405$ is the packing fraction of un-strained initial systems; $\varphi_{\rm f}=0.8516(5)$, $B=0.61(15)$ and $\tau=0.60(18)$ are fitting parameters.
+The error bar represents the standard error of the mean.
+Figure S6: Jamming fraction in three dimensions.
 (a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$.
-The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.648$.
+The intersection of curves gives $\varphi_{\rm J}^{\infty}={0.6479(2)}$.
 (b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm j}^{\infty})N^{1/2}$.
 The solid line represents the fitting according to Eq.
-(4), with two fitting parameters $u=0.077$ and $\sigma_{\varphi}=0.051$.
+(4), with two fitting parameters $u={0.077(2)}$ and $\sigma_{\varphi}={0.051(1)}$.
 Figure S7: Susceptibilities in three dimensions.
@@
 To draw these lines, we use $u=0.077$ and $\sigma_{\varphi}=0.051$ obtained from the fitting in Fig.
-S5b.
+S6b.
 Symbols in (a-c) have the same meanings as indicated in the legend of (a).